### 09 Mar 2026 v1

# Distributional Regression with Tabular Foundation Models: Evaluating Probabilistic Predictions via Proper Scoring Rules

Abstract
Prior-Data Fitted Networks (PFNs), such as TabPFN and TabICL, have revolutionized tabular deep learning by leveraging in-context learning for tabular data. These models are meant as foundation models for classification and regression settings and promise to greatly simplify deployment in practical settings because their performance is unprecedented (in terms of mean squared error or $R^{2}$, when measured on common benchmarks like TabArena or TALENT).
However, we see an important weakness of current benchmarks for the regression setting: the current benchmarks focus on evaluating win rates and performance using metrics like (root) mean squared error or $R^{2}$. Therefore, these leaderboards (implicitly and explicitly) push researchers to optimize for machine learning pipelines which elicit a good mean value estimate.
The main problem is that this approach only evaluates a point estimate (namely the mean estimator which is the Bayes estimator associated with the mean squared error loss). In this article we discuss the application of proper scoring rules for evaluating the goodness of probabilistic forecasts in distributional regression. We also propose to enhance common machine learning benchmarks with metrics for probabilistic regression. To improve the status quo and make the machine learning community aware of scoring rules for probabilistic regression, we advocate to use the continuous ranked probability score (CRPS) in benchmarks for probabilistic regression. However, we also illustrate that the choice of the scoring rule changes the inductive bias of the trained model (see section 4.1). We, therefore, advocate for finetuning or promptable tabular foundation models.
Keywords Fine-tuning $\cdot$ Benchmarking $\cdot$ Proper Scoring Rules $\cdot$ PFN
1 Introduction
The recent emergence of Prior-Data Fitted Networks (TabPFN) Hollmann et al. (2022, 2025) has shifted the paradigm of tabular learning from gradient boosted models to in-context learnersQu et al. (2026) which place attention on the right examples during the forward pass (without weight adaptation) of the underlying neural networks.
Besides the in-context learning, we recognize that both leading models TabPFN and also TabICLv2 perform forms of (discretized) distributional regression - an approach we greatly support since it allows for full probabilistic forecasts (in simple terms this means that the models allow to "predict a histogram over the support of potential outcomes", similar to the insets in figure 3).
[…]
A great development of the emergence of TabPFN is that it moves away from point estimates — such as the conditional median (e.g. arising from mean absolute error minimization) or the conditional mean (from minimizing e.g. mean squared error)—to full distributional regression. Joining the previous rage against the mean, we agree that point estimates may not offer the deepest insight into a problemKneib et al. (2023) since they fail to capture aleatoric uncertainty (uncertainty you see from one-data point to another one).
We show a simple toy model to illustrate problems with the usual ordinary least squares estimation which tries to elicit a mean estimate: a regression problem where the target variable has multiple modes (see figure 3 and the appendix for the description of the data generating process and the models involved) [^2] One might argue that such a bimodal response may be due to an unobserved confounder. However, we still need to work with the data at hand and, therefore, we argue that modeling the aleatoric uncertainty is important to make better decisions..
Figure 1: Toy-Model CRPS
Figure 2: Toy-Model CE
Figure 3: The images illustrate effects of training with CRPS or log-score (cross entropy loss) with same training data, same architecture but other loss function (CRPS vs log score) at finite amount of samples (1600 train samples, 400 test samples). For further description of the models and the data generating process see appendix.
One main takeaway from the toy model in figure 3 is that mean value estimates may fall in regions without realizations and may (depending on the utility) be completely useless (think about asking: "What could be the next outcome of a six sided die?" Even if you had a perfect mean value estimate telling you the mean value 3.5, then this would not even be a valid realization). Additionally, we also see that the usual construction of a prediction interval around the mean value estimate would cover a lot of empty space without samples (just by construction as an interval), which is inferior to estimating a conditional histogram like shown in the figure.
Interestingly, for the toy dataset, the Cross-entropy loss trained model in figure 3 differs significantly in terms of alignment with the ground truth data generating process, which indicates that at finite amount of samples:
  • the convergence speed during model training may depend on the chosen loss function - even if both loss functions derive from proper scoring rules
  • the sample efficiency of the complete training pipeline may depend on the loss function - even if both loss functions derive from proper scoring rules
In the table 11 in the appendix we observe that different models have different performance metrics, depending on the loss function even at otherwise identical training procedure. We also had observed this behavior on other openml datasets [^3] See https://github.com/PriorLabs/TabPFN/pull/689 for the pull request proposing CRPS for fine-tuning TabPFN.
A proper scoring rule $\mathbf{S}:\mathcal{F}\times\Omega\rightarrow\mathbb{R}$ guarantees that the expected score is minimized by the true distribution $Q\in\mathcal{F}$:
$\mathbb{E}_{Y\sim Q}[\mathbf{S}(Q,Y)]\leq\mathbb{E}_{Y\sim Q}[\mathbf{S}(F,Y)]\quad\forall F\in\mathcal{F}.$ (1)
However, in practice, we minimize the sample average score $\widehat{E[\mathbf{S}]}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{S}(F,y_{i})$. While $\widehat{E[\mathbf{S}]}$ is a consistent estimator of the expectation, the optimization path at finite $n$ is governed by the specific gradients $\nabla\mathbf{S}$ induced by the rule.
The Logarithmic Score, $\mathbf{S}_{Log}(F,y)=-\log f(y)$, exhibits extreme sensitivity to the tails of $\Omega$. Since $\nabla\mathbf{S}_{Log}\propto 1/f(y)$, the penalty for low-density observations grows toward infinity. At finite $n$, even a single outlier can dominate the gradient, forcing the optimizer to prioritize tail-fitting early, often at the expense of the distribution’s central mass.
Conversely, the Continuous Ranked Probability Score (CRPS) provides a more robust optimization landscape. This is best understood by expressing the CRPS as an integral over pinball losses $\rho_{\tau}$ for all quantiles $\tau\in[0,1]$:
$\text{CRPS}(F,y)=2\int_{0}^{1}\rho_{\tau}(F^{-1}(\tau),y)d\tau,$ (2)
where $\rho_{\tau}(u,y)=(y-u)(\tau-\mathbb{I}\{y<u\})$.
Because the gradient of the pinball loss is bounded (constant in magnitude), the CRPS gradient remains stable regardless of how far $y$ lies in the tails. Consequently, at finite $n$, CRPS-based estimation focuses on the global probability mass and is less susceptible to being pulled away by sparse observations in $\Omega$. While both rules are asymptotically equivalent in their target, their finite-sample "optimum" differs and also their behavior for tail-sensitivity (LogS) and robust mass-estimation (CRPS).
1.1 Important Quotes
  • "It is difficult to understand why statisticians commonly limit their inquiries to averages, and do not revel in more comprehensive views. Their souls seem as dull to the charm of variety as that of the native of one of our flat English counties, whose retrospect of Switzerland was that, if its mountains could be thrown into its lakes, two nuisances would be got rid of at once." (Galton, 1889) Kneib et al. (2023)
  • "Despite these well-argued perspectives, there has been little recognition that the common practice of requesting some point forecast, and then evaluating the forecasters by using some (set of) scoring function(s), is not a meaningful endeavor. In this article, we develop the perspectives of Murphy and Daan (1985) and Engelberg, Manski, and Williams (2009) and argue that effective point forecasting depends on guidance or directives, which can be given in one of two complementary ways, namely, by disclosing the scoring function ex ante to the forecaster, or by requesting a specific functional of the forecaster’s predictive distribution, such as the mean or a quantile." Gneiting (2011)
  • "Although every proper scoring rule can identify the true model or the conditional distribution as its minimum in the population sense, they exhibit different inductive biases or regularization effects in finite samples. Furthermore, the computational tractability varies substantially depending on the chosen scoring rule" Waghmare and Ziegel (2025)
1.2 Contributions
We greatly appreciate the recent improvement of tabular foundation models to move away from point estimates.
However, current regression benchmarks for Tabular Machine Learning like TabArena or TALENT are still stuck in evaluating "historically important" point estimates only.
We
  • propose the enhancement of regression benchmarks to account for probabilistic regression metrics derived from scoring rules, mainly CRPS by discussing a bimodal toy model.
  • we provide first evaluations of realTabPFNv2.5 and TabICLv2 in terms of proper scoring rules (CRPS and beta energy score).
  • we show how finetuning with different scoring rules than the one seen during pretraining in a finite-sample setting can enhance realTabPFNv2.5 performance (CRLS, beta energy score)
  • we show that in theory the choice of a scoring rule can have a significant impact in terms of the inductive bias of the model (having implications for finetuning and benchmarking)Merkle and Steyvers (2013); Waghmare and Ziegel (2025)
  • Additionally in the outlook, we start the discussion around a potential conflict between optimality of predictions with regard to user-supplied scoring rule and the goal to not adapt a general-purpose foundation model after pretraining.
2 Proposed Enhanced Reporting of Proper Scoring Rules in (Tabular) Regression Benchmarks
In the spirit of the seminal work by GneitingGneiting and Raftery (2007); Gneiting and Ranjan (2011); Gneiting (2011), we propose enhancing regression benchmarks for tabular foundation models by incorporating proper scoring rules. This approach shifts the focus from point-wise accuracy to a comprehensive evaluation of the predicted conditional probability distribution.
2.1 Log Score
Viewing discretized regression as a multi-class classification problem suggests the application of Cross-Entropy (CE), which corresponds to the log scoring rule which is local. This choice may be optimal if the final business objective is strictly to maximize the probability of the "correct" bin.
2.2 Continuous Ranked Probability Score (CRPS)
However, using cross entropy loss (meaning the log scoring rule) would ignore the distance between the prediction and the ground truth (bins are not even treated as having an ordinal order like in $1<2<3$ but more like completely independent concepts like "panda", "tiger", "fish"). To inject geometric awareness into the gradient, we proposed to use CRPS for finetuning TabPFN in December 2025444https://github.com/PriorLabs/TabPFN/pull/689 and saw CRPS being used in TabICLv2 training in February 2026Qu et al. (2026). In a regression context, predicting $y=100$ when the truth is $y=10$ should be penalized more heavily than predicting $y=11$. Still, standard Cross-Entropy treats both errors similarly if they fall into different wrong bins.
For a discretized regression output space, the discretized CRPS (equivalent to $\beta=1$ energy score) is given by:
$CRPS(F,y)\approx\sum_{i=1}^{N}(F(x_{i})-\mathbbm{1}_{\{x_{i}\geq y\}})^{2}\Delta x_{i},$ (3)
where $F(x_{i})$ is the predicted cumulative distribution function (e.g. obtained via cumsum of the softmax output of a neural network over the binned support). To perform optimum score estimation the expectation value (over all training samples) scoring rule needs to be taken and minimized.
However, CRPS weights all quantiles of the output space equally. This is likely, prohibitive in applications with asymmetric risk structure over the output space (see later discussion also for weighted CRPSGneiting and Ranjan (2011)).
We also think that at finite sample sizes the different scoring rules may differ (under otherwise constant conditions) in convergence speed and in sample efficiency (compare the results of the toy example in figure 3).
2.3 Continuous Ranked Logarithmic Scoring Rule (CRLS)
CRLS is defined by [^5] https://scoringrules.readthedocs.io/en/latest/theory.html
$\mathrm{CRLS}(F,y)=-\int_{-\infty}^{\infty}\log|F(x)+\mathbf{1}\{y\leq x\}-1|dx.$ (4)
2.4 Multivariate Enhancements
For $y\in\Omega\subseteq\mathbb{R}^{d}$ with $d>1$, univariate scores are insufficient for capturing dependence. We employ and suggest to benchmark for the Energy Score ($\mathbf{S}_{\beta}$) and Variogram Score ($\mathbf{S}_{VS,p}$) as also done in Landsgesell (2026).
The Energy Score for $\beta\in(0,2)$ is defined as:
$\mathbf{S}_{\beta}(F,y)=\mathbb{E}_{F}\|X-y\|^{\beta}-\frac{1}{2}\mathbb{E}_{F}\|X-X^{\prime}\|^{\beta}$ (5)
where $X,X^{\prime}\sim F$ are independent. For $d=1$ and $\beta=1$, $\mathbf{S}_{\beta}$ reduces to the CRPS.
The Variogram Score of order $p>0$ specifically targets the correlation structure by penalizing errors in the pairwise differences:
$\mathbf{S}_{VS,p}(F,y)=\sum_{j,k=1}^{d}w_{j,k}\left(|y_{j}-y_{k}|^{p}-\mathbb{E}_{F}|X_{j}-X_{k}|^{p}\right)^{2}$ (6)
where $w_{j,k}$ are weights. While $\mathbf{S}_{\beta}$ ensures strict propriety, $\mathbf{S}_{VS,p}$ provides superior sensitivity to misspecified dependencies between components of $y$.
3 Results
We adapt the finetuning code from github to start from the realTabPFNv2.5 checkpoint and implement custom loss functions (Beta energy score (with CRPS as special case), CRLS, and others).
We fit models on a diverse set of openML datasets for regression as shown in the various result tables. For each dataset we limit the amount of samples to randomly selected 3000 samples (random subsampling). We then perform 5 fold cross-validation on each subsampled dataset so that we can report performance on holdout data. On each split we evaluate a baseline model (always the realTabPFNv2.5) against another model (either a finetuned model or TabICLv2), so that we get paired performance metrics (per split). We report several performance metrics ($MAE$, RMSERMSE, $R^{2}$, $CRPS$, CRLSCRLS, $IntervalScore@95$). All metrics except R2 are better when they are lower. The improvement in $R^{2}$ is reported as absolute difference in $\Delta R^{2}*100$pp (to avoid statistical problems with reporting percent changes on fractions). The other improvements are reported as relative change from the baseline.
For simple interpretation we adopt the sign convention that a positive change is an improvement in the respective metric. For the metrics we first compute the change in the metric (at same split) and then compute the average afterwards.
Table 1: Finetuned TabPFN Regressor Hyperparameters
Category Hyperparameter Value
Ensemble n_estimators_finetune 1
Ensemble n_estimators_validation 8
Ensemble n_estimators_final_inference 8
Optimization learning_rate $1\times 10^{-5}$
Optimization weight_decay 0.1
Training epochs 600
Training early_stopping_patience 20
Data Handling n_finetune_ctx_plus_query_samples 20,000
Data Handling finetune_ctx_query_split_ratio 0.4
Loss Weights crps_loss_weight 0.0
Loss Weights crls_loss_weight 0.0
Loss Weights ce_loss_weight 0.0
Loss Weights mse_loss_weight 0.0
Loss Weights mae_loss_weight 0.0
Loss Weight custom implementation 1.0
Model Logic average_before_softmax True
3.1 Beta(1.8)-Energy Score Finetuning realTabPFNv2.5
We optimize with $\beta=1.8$ energy score and apply early stopping on MAE (changing the metric for early stopping showed strongly varying performance). We find improved performance in terms of $R^{2}$:
Table 2: Comparative Performance and Uncertainty Metrics (Finetuned vs. Base)
Metric Improvement (mean $\pm$ std) Median W/L/T
MAE +4.28% $\pm$ 10.06% +1.46% 61/21/8
RMSE +2.16% $\pm$ 4.39% +0.80% 61/24/5
$R^{2}$ +1.47pp $\pm$ 3.73pp +0.25pp 52/19/19
CRPS +2.76% $\pm$ 6.38% +0.79% 61/23/6
IntervalScore95 +2.24% $\pm$ 13.86% +0.94% 55/33/2
Table 3: Improvement by Dataset: Finetuned vs. Base
Dataset MAE Imp. RMSE Imp. $R^{2}$ Imp. CRPS Imp. IS95 Imp.
Abalone -0.03% +0.47% +0.39pp +0.78% +0.60%
Ailerons -0.04% -0.24% -0.06pp +0.18% -0.30%
Bike Sharing +1.49% +1.11% +0.13pp +1.08% -0.49%
Black Friday +0.16% +0.30% +0.30pp +0.25% -0.74%
Brazilian Houses -7.33% +3.43% +6.28pp +0.99% +5.64%
California Housing +0.91% +1.10% +0.48pp +1.25% +1.37%
Cpu Small +1.38% -0.12% -0.00pp -0.03% +0.92%
Diamonds +0.61% +1.26% +0.05pp +0.53% -1.46%
Digits +34.54% +6.29% +0.01pp +8.27% +13.74%
Elevators +0.02% +0.20% +0.03pp +0.21% +0.82%
House 16H -0.39% -0.06% -0.08pp -0.47% -0.69%
Kin8nm +1.83% +2.05% +0.30pp +2.01% +5.18%
Mercedes Benz +15.71% +6.94% +7.33pp +13.58% +7.27%
NYC Taxi +8.64% +4.73% +6.68pp +8.18% +3.50%
Protein +1.44% +0.26% +0.25pp +0.51% +1.12%
Pol +8.31% +3.30% +3.71pp +5.75% +9.80%
Puma8NH +7.87% +7.95% +0.77pp +6.85% +7.06%
Sulfur +1.85% -0.04% -0.02pp -0.17% -13.03%
Figure 4: Impact of beta=1.8 energy score finetuning on realTabPFNv2.5
3.2 CRLS-Finetuning realTabPFNv2.5
We implement the CRLS as a loss in tabpfn and use it here also as evaluation metric for early stopping.
Table 4: Overall Performance and Uncertainty Improvement (Finetuned vs. Base)
Metric Improvement (mean $\pm$ std) Median W/L/T
MAE +0.65% $\pm$ 10.13% +0.21% 54/41/5
RMSE +1.67% $\pm$ 4.60% +0.96% 73/25/2
$R^{2}$ +0.94pp $\pm$ 2.40pp +0.33pp 64/20/16
CRLS +2.27% $\pm$ 7.00% +1.47% 79/18/3
IntervalScore95 +3.87% $\pm$ 13.15% +1.76% 75/22/3
Table 5: Detailed Dataset Improvements: Finetuned vs. Base
Dataset MAE Imp. RMSE Imp. $R^{2}$ Imp. CRLS Imp. IS95 Imp.
Abalone -0.76% +0.42% +0.35pp +1.52% +2.58%
Ailerons +1.01% +0.65% +0.17pp +1.52% +0.86%
Ames Housing -0.21% +1.41% -0.23pp -0.75% -7.59%
Bike Sharing +1.86% +1.88% +0.21pp +1.19% +1.59%
Black Friday +0.23% +0.61% +0.61pp +1.09% +0.67%
Brazilian Houses -24.19% +3.67% +6.72pp -2.45% +23.84%
California Housing +0.92% +1.94% +0.74pp +2.48% +3.10%
Cpu Small +0.69% +0.05% +0.00pp +1.14% +1.45%
Diamonds +0.60% +1.27% +0.05pp +1.01% +1.15%
Digits +28.23% +7.48% -0.02pp +10.91% +3.09%
Elevators -0.54% +0.13% +0.02pp +0.34% +0.67%
House 16H -1.88% +0.42% +0.27pp +0.20% +0.61%
House Sales -0.71% -1.35% -0.34pp +0.22% -1.02%
Kin8nm -0.39% -0.77% -0.12pp -0.32% +0.97%
Mercedes Benz +3.46% +1.14% +1.22pp +2.12% +1.31%
NYC Taxi +4.71% +3.04% +3.90pp +4.19% +2.08%
Protein -1.47% +0.51% +0.45pp +1.43% +3.34%
Pol +7.37% +3.23% +3.64pp +8.15% +13.82%
Puma8NH +7.24% +7.41% +0.71pp +6.26% +6.37%
Sulfur -13.24% +0.36% +0.48pp +5.14% +18.50%
Figure 5: Impact of CRLS finetuning realTabPFNv2.5
3.3 CRLS performances of TabICLv2 versus realTabPFNv2.5
Table 6: Aggregate Performance and Uncertainty Improvement (TabICLv2 vs. realTabPFNv2.5)
Metric Improvement (mean $\pm$ std) Median W/L/T
MAE +1.41% $\pm$ 12.65% +0.16% 56/48/1
RMSE +1.85% $\pm$ 10.02% +0.96% 65/34/6
$R^{2}$ +3.22pp $\pm$ 17.03pp +0.17pp 56/27/22
CRLS +6.01% $\pm$ 10.95% +3.49% 84/19/2
IntervalScore95 +5.14% $\pm$ 14.18% +2.15% 70/35/0
Table 7: Dataset-Specific Improvement Metrics
Dataset MAE Imp. RMSE Imp. $R^{2}$ Imp. CRLS Imp. IS95 Imp.
Abalone +0.22% +0.92% +0.76pp +3.37% +4.94%
Ailerons -1.69% -1.01% -0.29pp +1.55% +0.23%
Ames Housing +0.37% +8.13% +32.38pp +4.95% -13.51%
Bike Sharing +0.21% -0.62% -0.06pp +2.83% -0.13%
Black Friday +0.59% +1.17% +1.18pp +0.34% -2.20%
Brazilian Houses -13.82% +3.41% +6.13pp +25.02% +31.86%
California Housing +4.87% +5.25% +1.91pp +8.33% +6.34%
Cpu Small +1.49% +0.08% +0.00pp +6.17% +3.52%
Diamonds +1.15% +1.62% +0.06pp +4.33% +4.93%
Digits +42.92% +21.15% +0.89pp +34.26% +43.52%
Elevators -0.20% +0.57% +0.10pp +1.55% -0.55%
House 16H +0.76% +2.92% +1.92pp +5.30% +5.42%
House Sales -5.00% -10.86% -2.37pp +0.11% -3.01%
Kin8nm +0.59% -0.29% -0.04pp +1.52% +3.26%
Mercedes Benz +18.83% +9.57% +9.92pp +20.31% +13.27%
NYC Taxi -7.26% +3.99% +9.25pp -5.47% -6.48%
Protein +2.63% +1.71% +1.55pp +4.28% +3.15%
Pol +10.39% +4.05% +4.56pp +11.22% +7.74%
Puma8NH -13.29% -13.86% -1.50pp -13.66% -10.89%
Sulfur -10.13% +1.27% +1.61pp +6.38% +16.39%
Visualizing Soil -4.08% -0.36% -0.32pp +3.50% +0.08%
Figure 6: Impact of changing from RealTabPFNv2.5 to TabICLv2: Even stronger outperformance of TabICLv2 over realTabPFNv2.5 on CRLS than seen before by finetuning realTabPFNv2.5 on CRLS
3.4 CRPS performances of realTabPFNv2.5 and TabICLv2
Figure 7: Performance Improvement of TabICLv2 over realTabPFNv2.5 on CRPS (and other metrics)
Table 8: Overall Model Improvement (TabICLv2 vs. realTabPFNv2.5)
Metric Improvement (mean $\pm$ std) Median W/L/T
MAE +1.41% $\pm$ 12.65% +0.16% 56/48/1
RMSE +1.85% $\pm$ 10.02% +0.96% 65/34/6
$R^{2}$ +3.22pp $\pm$ 17.03pp +0.17pp 56/27/22
CRPS +2.01% $\pm$ 10.17% +0.73% 59/45/1
IntervalScore95 +5.14% $\pm$ 14.18% +2.15% 70/35/0
Table 9: Dataset-Specific Performance Gains
Dataset MAE Imp. RMSE Imp. $R^{2}$ Imp. CRPS Imp. IS95 Imp.
Abalone +0.22% +0.92% +0.76pp +0.86% +4.94%
Ailerons -1.69% -1.01% -0.29pp -2.20% +0.23%
Ames Housing +0.37% +8.13% +32.38pp -1.81% -13.51%
Bike Sharing +0.21% -0.62% -0.06pp -0.75% -0.13%
Black Friday +0.59% +1.17% +1.18pp +0.61% -2.20%
Brazilian Houses -13.82% +3.41% +6.13pp +3.63% +31.86%
California Housing +4.87% +5.25% +1.91pp +5.84% +6.34%
Cpu Small +1.49% +0.08% +0.00pp +1.82% +3.52%
Diamonds +1.15% +1.62% +0.06pp +1.45% +4.93%
Digits +42.92% +21.15% +0.89pp +28.19% +43.52%
Elevators -0.20% +0.57% +0.10pp -0.17% -0.55%
House 16H +0.76% +2.92% +1.92pp +1.82% +5.42%
House Sales -5.00% -10.86% -2.37pp -4.79% -3.01%
Kin8nm +0.59% -0.29% -0.04pp +0.20% +3.26%
Mercedes Benz +18.83% +9.57% +9.92pp +20.91% +13.27%
NYC Taxi -7.26% +3.99% +9.25pp -8.96% -6.48%
Protein +2.63% +1.71% +1.55pp +2.79% +3.15%
Pol +10.39% +4.05% +4.56pp +6.73% +7.74%
Puma8NH -13.29% -13.86% -1.50pp -15.15% -10.89%
Sulfur -10.13% +1.27% +1.61pp +1.78% +16.39%
Visualizing Soil -4.08% -0.36% -0.32pp -0.56% +0.08%
4 Choice of Scoring Rules for Decision Making
The choice of a scoring rule matters because different scoring rules have different inductive biases Waghmare and Ziegel (2025); Merkle and Steyvers (2013). This is the reason why the scoring rule used for solving a business problem should be chosen carefully:
  • Oesterheld and Conitzer (2020) describes "Decision Scoring Rules"
  • Winkler et al. describe tailored scoring rules for decision making from a utility perspectiveJohnstone et al. (2011).
  • The very insightful introduction example in Johnstone et al. (2011), page 2, is about making "useful" probabilistic weather forecasts for a specific use-case (an orchard) and we drew a lot of inspiration reading it - this example is very intuitive we suggest to read itJohnstone et al. (2011).
  • Merkle et al.Merkle and Steyvers (2013) in their paper "Choosing a strictly proper scoring rule" describe an important effect: The "optimal forecaster" depends on the choice of a scoring rule (while all scoring rules are individually minimized by a forecast which is equal to the true distribution).
In other words, different scoring rules penalize different types of errors differently and thus the optimal forecast (the one which minimizes the expected score) can differ based on the scoring rule used. Merkle already points this out and finds that different forecasters are ranked very differently by different strictly proper scoring rules. They for example find in table 1 that the spearman correlation for the ranking of forecasters between the brier score and the log score is only 0.15. They then go on and investigate a parametrizable family of scoring rules and see how the ranking generally depends on the choice of the parameters.
For this article we follow this idea and port it to proper scoring rules which can be used in the regression setting. We choose the $\beta$-energy score family and show that the optimal model depends on the choice of $\beta$. To show this and for keeping the arguments as simple as possible we consider
4.1 Invesigating the effect of the choice of scoring rule on the optimal forecast
For simplicity [^6] While the following argument describes concentrated probabilistic forecasts, the argument remains untouched even if we move to not so extremely concentrated forecasts. of the argument, we consider a probabilistic forecast of the following kind:
$\hat{p}_{i}(Y=y|X=x)=\delta(y-f_{i}(x)),$ (7)
where $i\in\{A,B,C,D,E\}$ indexes different forecasting models (e.g. from different checkpoints of a training) and $\delta()$ is the Dirac delta distribution (which collapses the probability distribution to a "pointy" probability density). In this toy model (again for simplicity of the argument) we impose the ground truth
$p_{g}(Y=y|X=x)=\delta(y-g(x)).$ (8)
We choose $x\in[-3,3]$ (np.linspace(-3,3,1000)).
  • $g(x)=\sin(x)$
  • Model A: Small consistent bias, $f_{A}(x)=g(x)+0.1$
  • Model B: Perfect with a large localized deviation, $f_{B}(x)=g(x)$ with an additional $2.0$ added to the middle 50 points
  • Model C: Scaled version of the true function, $f_{C}(x)=0.8\cdot g(x)$
  • Model D: Shifted version of the true function, $f_{D}(x)=\sin(x+0.2)$
  • Model E: Simple baseline at zero, $f_{E}(x)=0$
A typical user would now wonder which model is best? To get an intuition, we look at the models in Figure 8.
Figure 8: Toy Models for investigating the effect of the choice of scoring rule on the optimal forecast
Obviously the different models make different erros which can be seen in the absolute error plot in Figure 9. Some models make small errors everywhere (model A), some models make large errors in a small region (model B), some models make medium sized errors everywhere (model C) and some models make small errors everywhere but are shifted (model D).
Figure 9: Absolute errors for the toy models
Now we can apply different scoring rules (from the $\beta$-energy family) to see which model is optimal for which scoring rule. For a fixed choice of $\beta$ (one point on the x-axis in Figrue 10) the better model has the lower beta energy score.
Figure 10: Decision based on scoring rules for the toy models
We observe that while the model B is the best ranked model for each $\beta\lesssim 1$ for $1\lesssim\beta$ this changes and also other models switch rank at other values of $\beta$, see figure 11.
Figure 11: Ranking of toy models based on scoring rules
Fundamentally this behavior arises from different scoring rules rating the severity of errors in Figure 9 differently (and there is even scoring rules wich penalize errors depending on the direction of the error, see for example asymmetric CRPSIacopini et al. (2023) which may be important e.g. in finance where errors have different severity depending on the direction of the error: making 100 percent more money is less severe than loosing 100 percent of money).
For simple probabilistic point masses like shown here we can even make an analytical statement: The beta energy score with $\beta=1$ favors minimization of the mean absolute error (i.e. prefers estimating a median estimate) while for $\beta=2$ the scoring rule prefers a minimization of mean squared error.
Derivation.
The $\beta$-energy score for a forecast distribution $P$ and observation $y\in\mathbb{R}$ is defined as
$\mathrm{ES}_{\beta}(P,\,y)=\mathbb{E}_{X\sim P}\bigl[|X-y|^{\beta}\bigr]-\tfrac{1}{2}\,\mathbb{E}_{X,\,X^{\prime}\sim P}\bigl[|X-X^{\prime}|^{\beta}\bigr],\qquad 0<\beta\leq 2,$ (9)
where $X$ and $X^{\prime}$ are independent draws from $P$.
Now suppose the forecast is a probabilistic point mass, i.e. $P=\delta_{\hat{m}}$ for some point estimate $\hat{m}\in\mathbb{R}$. Both $X$ and $X^{\prime}$ equal $\hat{m}$ almost surely, so the second term in (9) vanishes and the score reduces to
$\mathrm{ES}_{\beta}(\delta_{\hat{m}},\,y)=|\hat{m}-y|^{\beta}.$ (10)
The optimal point estimate $\hat{m}^{*}$ minimises the expected score over observations $y\sim Q$ for some true data distribution $Q$:
$\hat{m}^{*}=\operatorname*{arg\,min}_{\hat{m}}\;\mathbb{E}_{y\sim Q}\!\left[|\hat{m}-y|^{\beta}\right].$ (11)
Case $\beta=1$.  Equation (11) becomes $\min_{\hat{m}}\,\mathbb{E}[|\hat{m}-y|]$, i.e. minimisation of the mean absolute error (MAE). The unique minimiser is the median of $Q$, since
$\frac{\partial}{\partial\hat{m}}\,\mathbb{E}[|\hat{m}-y|]=\Pr(y<\hat{m})-\Pr(y>\hat{m})=0\;\Longleftrightarrow\;\hat{m}=\mathrm{median}(Q).$ (12)
Case $\beta=2$.  Equation (11) becomes $\min_{\hat{m}}\,\mathbb{E}[(\hat{m}-y)^{2}]$, i.e. minimisation of the mean squared error (MSE). Setting the derivative to zero gives
$\frac{\partial}{\partial\hat{m}}\,\mathbb{E}\!\left[(\hat{m}-y)^{2}\right]=2\,(\hat{m}-\mathbb{E}[y])=0\;\Longleftrightarrow\;\hat{m}^{*}=\mathbb{E}[y]=\mathrm{mean}(Q).$ (13)
Hence the $\beta$-energy score with $\beta=2$ is minimised by reporting the mean of the predictive distribution.
The choice of the scoring rule determines how different types of errors are penalized and thus the optimal forecast (the one which minimizes the expected score) can differ based on the scoring rule used. This is true, despite every scoring rule being strictly proper (i.e. every scoring rule is minimized by the true distribution). The reason is that different scoring rules penalize different types of errors differently and thus the optimal forecast (the one which minimizes the expected score) can differ based on the scoring rule used.
We leave it as an open question how to best choose a scoring rule for general purpose tabular foundation model pretraining and how to enable adaptation to arbitrary business case with high severity for certain errors (like in finance or pharmaceutical production). We describe some possible routes to explore in the outlook section 7.
5 Discussion
Across the datasets evaluated, finetuning with beta-energy score ($\beta=1.8$) and CRLS both tend to improve realTabPFNv2.5 on the majority of metrics and datasets (as indicated by the W/L/T counts in Tables 2 and 4). However, improvements are modest in median terms (typically $2\%$) and vary considerably across datasets.
The aggregate results further suggest that TabICLv2 may be a stronger probabilistic model than realTabPFNv2.5 in terms of CRPS, CRLS, and Interval Score, with TabICLv2 winning on the majority of evaluated datasets (Tables 6, 8). The margin varies substantially across datasets (e.g. large gains on Mercedes Benz and Pol, but losses on Puma8NH and House Sales).
Additional experiments examining how finetuning performance scales with dataset size would be valuable to better understand whether the observed differences are robust across data regimes.
We illustrate in section 4.1 that the choice of scoring rule can change the ranking of forecasting models — at least in the considered toy setting of Dirac-delta forecasts over a smooth ground truth function. Under this simplification, the result follows analytically from the $\beta$-energy score reducing to an $L^{\beta}$ norm. Whether this ranking sensitivity carries over to actual trained distributional models in realistic tabular settings remains an open empirical question. That said, the theoretical argument and the empirical literature Waghmare and Ziegel (2025); Merkle and Steyvers (2013) both support the view that the choice of scoring rule for pretraining or finetuning a tabular foundation model is non-trivial and may affect which use-cases the model is best suited for Oesterheld and Conitzer (2020); Johnstone et al. (2011).
6 Limitations: Rare or Extreme Events
We want to highlight that optimal scoring rule estimation (minimizing/maximizing the expected value of a scoring rule) is inherently not well suited for eliciting properties which are only realized in rare events (fall in the area where no data was observed or in 1D are in the tail of a distribution). This was well argued in recent years for the case of one dimensional distributions but equally carries over to higher dimensions. Lerch et al. (2017); Brehmer and Strokorb (2019) Inherently, this problem is related to epistemic uncertainty (uncertainty due to not having training data). Based on this insight also PFNs (trained to optimize mean scoring rules) should be applied with care if risk associated mainly with rare (tail) events or events outside the support of the examples in the supplied context.
7 Outlook
While we discussed how to properly measure the goodness of a probabilistic forecast - namely by applying proper scoring rules, we highlight, that this discussion is not finished. Gneiting pointed out in Gneiting and Ranjan (2011) that situations with asymmetric risk (finance, weather forecasting etc.) may require to optimize for specific point estimates (via scoring functions suitable to elicit a specific point estimate). In such a case, we wonder how tabular foundation models would adapt to specific user supplied scoring functions (so that optimality is guaranteed). We point out a conflict of
  • a foundation model which is not altered by a user at all after pretraining is finished
  • the need to take into account that the optimal scoring function depends on the use-case at hand (must be supplied by the user since the user intention dictates the utility).
We see that Fine-Tuning of foundation models could be one route to enable optimal predictions for arbitrary scoring functions. Another route could be to enable task-token conditional probability prediction in case of asymmetric risk structures (a user would prompt the model with a desired scoring rule which should be optimal, but this behaviour would need to be pretrained). Such a task-token conditional probability prediction could output probability estimates which are optimal under some parametrizable family of scoring rules (like e.g. the beta energy score family). However there is infinitely many scoring rules and different families of scoring rules so that it is unlcear if pre-training with taks-tokens can in general adapt to arbitrary scoring rules as objectives. Potentially, tabular foundation models for high stakes applications need to make sensible choices for task conditioning for certain applications (finance, pharma production, etc.).
7.1 Example: Weighted CRPS (wCRPS)
Accounting for asymmetric risk structures, a modeler can choose to apply the different versions of weighted CRPSGneiting and Ranjan (2011) - a family of strictly proper scoring rules which enables e.g. emphasizing prediction quality for the left/right tail of a distribution which makes sense when dealing with e.g. energy forecasting or forecasting weather events with asymmetric cost.
$L_{wCRPS}\approx\sum_{i=1}^{N}w(x_{i})(F(x_{i})-\mathbbm{1}_{\{x_{i}\geq y\}})^{2}\Delta x_{i},$ (14)
7.2 Example: Interval Score
The Interval ScoreGneiting and Raftery (2007) targets the $(1-\alpha)$ central interval and is strictly proper, directly answering "how well is the interval estimated?" for a given significance level $\alpha$:
$S_{\alpha}(F,y)=(u-l)+\frac{2}{\alpha}(l-y)\mathbb{I}\{y<l\}+\frac{2}{\alpha}(y-u)\mathbb{I}\{y>u\}$ (15)
where $[l,u]$ are the $\alpha/2$ and $1-\alpha/2$ quantiles of the prediction $F$.
This specific scoring rule highlights that different scoring rules favor optimization in distinct regions of the probability mass function (PMF); in the extreme case of the interval score, the two boundary quantiles are of primary importance. Because we hypothesize that each scoring rule aligns with a specific utility, general-purpose models may fail to achieve optimality for specialized downstream tasks where the cost of error is non-uniform across the domain $\Omega$.
8 Acknowledgement
We thank David Holzmüller for a very engaging and supportive email exchange in March 2026. We thank Noah Hollmann for a very polite and engaging call just a few days prior to Christmas 2025.
9 Appendix
9.1 Toy-Models to investigate finite sample effects
9.1.1 Data Generating Process
We simulate a dataset where the target variable $y$ exhibits a bimodal conditional distribution given a feature $x$. This is achieved by generating $y$ values from two distinct linear relationships:
$y_{1}=m_{1}x+c_{1}+\epsilon_{1}$ (16)
$y_{2}=m_{2}x+c_{2}+\epsilon_{2}$ (17)
where $\epsilon_{1},\epsilon_{2}\sim N(0,\sigma^{2})$ represent independent Gaussian noise. For each observation of $x$, the corresponding $y$ is randomly sampled from either $y_{1}$ or $y_{2}$, creating a bimodal output distribution $P(y|x)$.
9.1.2 Toy Models
To evaluate the efficacy of probabilistic forecasting in such a scenario, we compare several modeling strategies:
Baseline Regression Models
  • Least Squares (LS): A Random Forest Regressor trained with a squared error criterion, approximating the conditional mean: $E[y|x]$.
  • Least Absolute Deviations (LAD): A Random Forest Regressor trained with an absolute error criterion, approximating the conditional median: $\text{Median}(y|x)$.
Probabilistic Classification Models These models predict probabilities across $N=50$ discrete bins of the $y$-variable. The neural networks are optimized using the Adam optimizer and are simple multi-layer perceptron (MLP) featuring an input layer, two hidden layers with 64 neurons each using $\tanh$ activations, and a Softmax output layer for $n$ bins.
Hyperparameter PyTorch CRPS Model PyTorch Cross-Entropy Model
Network Architecture
Input Dimension 1 1
Hidden Layer 1 Neurons 64 64
Hidden Layer 2 Neurons 64 64
Output Layer Neurons 50 (num_bins) 50 (num_bins)
Activation Function Tanh (hidden), Softmax (output) Tanh (hidden), Softmax (output)
Training Parameters
Loss Function RPS Loss CrossEntropyLoss
Optimizer Adam Adam
Learning Rate 0.001 0.001
Number of Epochs 100 100
Number of Bins (num_bins) 50 50
Table 10: Neural Network Hyperparameter Summary
  • PyTorch Cross-Entropy (CE) Model: A feed-forward neural network trained with Cross-Entropy loss (log score), aiming to maximize the probability of the true bin.
  • PyTorch Ranked Probability Score (RPS) Model: A neural network trained directly to minimize the expected CRPS, which optimizes the accuracy of the cumulative distribution function (CDF).
  • XGBoost RPS Model: An XGBoost vectorleaf classifier utilizing a custom objective function that directly optimizes the Continuous Ranked Probability Score
    1import numpy as np
    2import torch
    3import torch.nn.functional as F
    4
    5def crps_objective(preds, dtrain):
    6 """
    7 Improved RPS Objective for XGBoost.
    8 Optimized for multi-class flattening and numerical stability.
    9 """
    10 # 1. Labels
    11 labels = dtrain.get_label().astype(int)
    12
    13 # 2. Reshape preds if multi-class (XGBoost passes them flat)
    14 num_samples = labels.shape[0]
    15 num_bins = preds.size // num_samples
    16 preds = preds.reshape(num_samples, num_bins)
    17
    18 # 3. Softmax for Probabilities (using Torch for speed/consistency)
    19 preds_tensor = torch.tensor(preds, dtype=torch.float32)
    20 probs = F.softmax(preds_tensor, dim=1).numpy()
    21
    22 # 4. One-Hot Targets
    23 target_one_hot = np.zeros_like(probs)
    24 target_one_hot[np.arange(num_samples), labels] = 1
    25
    26 # 5. CDFs
    27 pred_cdf = np.cumsum(probs, axis=1)
    28 target_cdf = np.cumsum(target_one_hot, axis=1)
    29 cdf_diff = pred_cdf - target_cdf # (B, K)
    30
    31 # 6. Gradient Calculation
[…]
    33 grad_p = 2 * np.flip(np.cumsum(np.flip(cdf_diff, axis=1), axis=1), axis=1)
    34
    35 # Convert grad w.r.t probability to grad w.r.t logit (z_j)
    36 # grad_z_j = prob_j * (grad_p_j - sum_i(grad_p_i * prob_i))
    37 weighted_grad_p = np.sum(grad_p * probs, axis=1, keepdims=True)
    38 grad_z = probs * (grad_p - weighted_grad_p)
    39
    40 # 7. Hessian Calculation
    41 # Robust approximation: 2 * prob * (1 - prob)
    42 hess_z = 2.0 * probs * (1.0 - probs)
    43
    44 return grad_z.flatten(), hess_z.flatten()
    45
    46def rps_metric(preds, dtrain):
    47 """Evaluation metric for RPS."""
    48 labels = dtrain.get_label().astype(int)
    49 num_samples = labels.shape[0]
    50 num_bins = preds.size // num_samples
    51
    52 preds = preds.reshape(num_samples, num_bins)
    53 probs = F.softmax(torch.tensor(preds), dim=1).numpy()
    54
    55 target_one_hot = np.zeros_like(probs)
    56 target_one_hot[np.arange(num_samples), labels] = 1
    57
    58 pred_cdf = np.cumsum(probs, axis=1)
    59 target_cdf = np.cumsum(target_one_hot, axis=1)
    60
    61 rps = np.mean(np.sum((pred_cdf - target_cdf)**2, axis=1))
    62 return ’RPS’, rps
    63
    64import xgboost as xgb
    65
    66# 1. Convert training data
    67dtrain = xgb.DMatrix(X_train, label=y_train_labels)
    68
    69# 2. Parameters
    70params = {
    71 ’objective’: ’rank:pairwise’,
    72 ’eval_metric’: ’rmse’,
    73 ’tree_method’: ’hist’,
    74 ’multi_strategy’: ’multi_output_tree’,
    75 ’num_class’: num_bins,
    76 ’eta’: 0.1,
    77 ’max_depth’: 3,
    78 ’subsample’: 0.8,
    79 ’colsample_bytree’: 0.8,
    80 ’min_child_weight’: 5,
    81 ’lambda’: 1,
    82 ’alpha’: 0.5,
    83}
    84
    85# 3. Train model
    86num_boost_round = 100
    87xgb_rps_model = xgb.train(
    88 params,
    89 dtrain,
    90 num_boost_round=num_boost_round,
    91 obj=crps_objective,
    92)
9.1.3 Model Performance Comparison
The performance is evaluated using Mean Squared Error (MSE), Mean Absolute Error (MAE), and the Ranked Probability Score (CRPS).
Table 11: Performance Metrics Comparison
Model MSE $\downarrow$ MAE $\downarrow$ CRPS $\downarrow$
Least Squares (Mean) 44.0942 4.9899 8.1750
Least Absolute Deviations (Median) 43.4364 4.9779 8.1600
PyTorch CRPS Model 31.8049 4.5400 4.5798
PyTorch Cross-Entropy Model 37.5715 5.2336 6.1814
XGBoost CRPS Model 37.8305 4.4383 4.6355
9.1.4 Summary
In this bimodal dataset scenario:
  • Regression Failure: Least Squares and Least Absolute Deviations models struggle to capture the underlying distribution, as point estimates often fall in the low-probability "valley" between modes.
  • RPS Advantage: Both the PyTorch and XGBoost models trained with the custom CRPS objective provide superior probabilistic forecasts.
  • Conclusion: While MSE and MAE (as specific scoring functions, see Gneiting and Ranjan (2011)) provide insights into point prediction, CRPS is the appropriate metric for evaluating complex conditional distributions. The XGBoost model with the custom CRPS objective demonstrated the most robust performance.

### 13 Mar 2026 v2

@@
 Abstract
-Prior-Data Fitted Networks (PFNs), such as TabPFN and TabICL, have revolutionized tabular deep learning by leveraging in-context learning for tabular data.
-These models are meant as foundation models for classification and regression settings and promise to greatly simplify deployment in practical settings because their performance is unprecedented (in terms of mean squared error or $R^{2}$, when measured on common benchmarks like TabArena or TALENT).
-However, we see an important weakness of current benchmarks for the regression setting: the current benchmarks focus on evaluating win rates and performance using metrics like (root) mean squared error or $R^{2}$.
-Therefore, these leaderboards (implicitly and explicitly) push researchers to optimize for machine learning pipelines which elicit a good mean value estimate.
-The main problem is that this approach only evaluates a point estimate (namely the mean estimator which is the Bayes estimator associated with the mean squared error loss).
-In this article we discuss the application of proper scoring rules for evaluating the goodness of probabilistic forecasts in distributional regression.
-We also propose to enhance common machine learning benchmarks with metrics for probabilistic regression.
-To improve the status quo and make the machine learning community aware of scoring rules for probabilistic regression, we advocate to use the continuous ranked probability score (CRPS) in benchmarks for probabilistic regression.
-However, we also illustrate that the choice of the scoring rule changes the inductive bias of the trained model (see section 4.1).
-We, therefore, advocate for finetuning or promptable tabular foundation models.
+Tabular foundation models such as TabPFN and TabICL already produce full predictive distributions, yet the benchmarks used to evaluate them (TabArena, TALENT, and others) still rely almost exclusively on point-estimate metrics (RMSE, $R^{2}$).
+This mismatch implicitly rewards models that elicit a good conditional mean while ignoring the quality of the predicted distribution.
+We make two contributions.
+First, we propose supplementing standard point metrics with proper scoring rules (CRPS, CRLS, and the Interval Score) and provide a head-to-head comparison of realTabPFNv2.5 and TabICLv2 with regards to some proper scoring rules across 20 OpenML regression datasets.
+Second, we show analytically and empirically that different proper scoring rules induce different model rankings and different inductive biases during training, even though each rule is individually minimized by the true distribution.
+Fine-tuning realTabPFNv2.5 with scoring rules not seen during pretraining (CRLS, $\beta{=}1.8$ energy score) yields consistent improvements on the corresponding metrics, confirming that the training loss shapes the model beyond what propriety alone guarantees.
+Together, these findings argue for (i) reporting distributional metrics in tabular regression benchmarks and (ii) making the training objective of foundation models adaptable (via fine-tuning or task-token conditioning) to the scoring rule relevant to the downstream decision problem.
 Keywords Fine-tuning $\cdot$ Benchmarking $\cdot$ Proper Scoring Rules $\cdot$ PFN
@@
 (2026) which place attention on the right examples during the forward pass (without weight adaptation) of the underlying neural networks.
-Besides the in-context learning, we recognize that both leading models TabPFN and also TabICLv2 perform forms of (discretized) distributional regression - an approach we greatly support since it allows for full probabilistic forecasts (in simple terms this means that the models allow to "predict a histogram over the support of potential outcomes", similar to the insets in figure 3).
+Both leading models, TabPFN and TabICLv2, perform forms of discretized distributional regression, an approach we strongly support, as it enables full probabilistic forecasts (i.e., the models can predict a histogram over the support of potential outcomes, as illustrated by the insets in Figure 3).
 […]
-A great development of the emergence of TabPFN is that it moves away from point estimates — such as the conditional median (e.g.
-arising from mean absolute error minimization) or the conditional mean (from minimizing e.g.
-mean squared error)—to full distributional regression.
-Joining the previous rage against the mean, we agree that point estimates may not offer the deepest insight into a problemKneib et al.
-(2023) since they fail to capture aleatoric uncertainty (uncertainty you see from one-data point to another one).
-We show a simple toy model to illustrate problems with the usual ordinary least squares estimation which tries to elicit a mean estimate: a regression problem where the target variable has multiple modes (see figure 3 and the appendix for the description of the data generating process and the models involved) [^2] One might argue that such a bimodal response may be due to an unobserved confounder.
+A notable contribution of TabPFN is its shift away from point estimates (such as the conditional median arising from MAE minimization, or the conditional mean from MSE minimization) toward full distributional regression.
+We agree that point estimates often fail to offer the deepest insight into a problem Kneib et al.
+(2023), as they cannot capture aleatoric uncertainty, i.e., the inherent variability from one data point to another.
+As Galton already noted in 1889:
+“It is difficult to understand why statisticians commonly limit their inquiries to averages, and do not revel in more comprehensive views.
+Their souls seem as dull to the charm of variety as that of the native of one of our flat English counties, whose retrospect of Switzerland was that, if its mountains could be thrown into its lakes, two nuisances would be got rid of at once.” Kneib et al.
+(2023)
+Recent work has explored a range of strategies for adapting tabular foundation models, including meta-learning, supervised fine-tuning, and parameter-efficient fine-tuning (PEFT) Tanna et al.
+(2026, 2025).
+However, to our knowledge, these studies focus on whether and how to fine-tune (in terms of parameter update strategies), while the inductive bias introduced by the choice of loss function (scoring rule) for a probabilistic forecast has not been systematically investigated.
+1.1 Contributions
+Current regression benchmarks for tabular machine learning (e.g.
+TabArena, TALENT) evaluate models exclusively through point-estimate metrics such as RMSE or $R^{2}$.
+This practice implicitly rewards models that elicit a good conditional mean while ignoring the quality of the full predictive distribution, a critical shortcoming given that leading tabular foundation models (TabPFN, TabICLv2) already produce distributional outputs.
+Our contributions are twofold:
+(I) Towards probabilistic regression benchmarks.
+  • We argue, through a bimodal toy model and 20 OpenML regression datasets, that point-estimate leaderboards are insufficient and can be misleading when the predictive distribution is multimodal or heavy-tailed.
+  • We propose a concrete benchmark enhancement that supplements MAE, RMSE, and $R^{2}$ with distributional metrics (CRPS, CRLS, and Interval Score, $\mathrm{IS}_{95}$) and provide a head-to-head evaluation of realTabPFNv2.5 and TabICLv2 using proper scoring rules.
+  • We additionally outline multivariate extensions (Energy Score, Variogram Score) for benchmarking future joint predictive distributions.
+(II) Scoring-rule inductive bias and the need for adaptable training objectives.
+  • We show analytically (via the $\beta$-energy score family) and empirically that different proper scoring rules induce different model rankings, even though each rule is individually minimized by the true distribution Merkle and Steyvers (2013); Waghmare and Ziegel (2025).
+This may imply that no single pretraining loss can be universally optimal across all downstream evaluation criteria - it is essentially the question of how different errors are penalized and a global model may not be optimal for every application.
+  • We demonstrate that fine-tuning realTabPFNv2.5 with scoring rules not seen during pretraining (CRLS, $\beta{=}1.8$ energy score) yields consistent improvements over the base model on the corresponding metrics across the majority of the 20 benchmark datasets, confirming that the inductive bias of the training loss can be shifted post hoc.
+2 Motivating Example: Why Point Estimates Fail
+We illustrate the limitations of point-estimate evaluation with a simple toy model: a regression problem where the target variable has multiple modes (see Figure 3 and the appendix for the description of the data generating process and the models involved) [^2] One might argue that such a bimodal response may be due to an unobserved confounder.
 However, we still need to work with the data at hand and, therefore, we argue that modeling the aleatoric uncertainty is important to make better decisions..
@@
 Figure 2: Toy-Model CE
-Figure 3: The images illustrate effects of training with CRPS or log-score (cross entropy loss) with same training data, same architecture but other loss function (CRPS vs log score) at finite amount of samples (1600 train samples, 400 test samples).
-For further description of the models and the data generating process see appendix.
-One main takeaway from the toy model in figure 3 is that mean value estimates may fall in regions without realizations and may (depending on the utility) be completely useless (think about asking: "What could be the next outcome of a six sided die?" Even if you had a perfect mean value estimate telling you the mean value 3.5, then this would not even be a valid realization).
-Additionally, we also see that the usual construction of a prediction interval around the mean value estimate would cover a lot of empty space without samples (just by construction as an interval), which is inferior to estimating a conditional histogram like shown in the figure.
-Interestingly, for the toy dataset, the Cross-entropy loss trained model in figure 3 differs significantly in terms of alignment with the ground truth data generating process, which indicates that at finite amount of samples:
-  • the convergence speed during model training may depend on the chosen loss function - even if both loss functions derive from proper scoring rules
-  • the sample efficiency of the complete training pipeline may depend on the loss function - even if both loss functions derive from proper scoring rules
-In the table 11 in the appendix we observe that different models have different performance metrics, depending on the loss function even at otherwise identical training procedure.
-We also had observed this behavior on other openml datasets [^3] See https://github.com/PriorLabs/TabPFN/pull/689 for the pull request proposing CRPS for fine-tuning TabPFN.
-A proper scoring rule $\mathbf{S}:\mathcal{F}\times\Omega\rightarrow\mathbb{R}$ guarantees that the expected score is minimized by the true distribution $Q\in\mathcal{F}$:
-$\mathbb{E}_{Y\sim Q}[\mathbf{S}(Q,Y)]\leq\mathbb{E}_{Y\sim Q}[\mathbf{S}(F,Y)]\quad\forall F\in\mathcal{F}.$ (1)
-However, in practice, we minimize the sample average score $\widehat{E[\mathbf{S}]}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{S}(F,y_{i})$.
-While $\widehat{E[\mathbf{S}]}$ is a consistent estimator of the expectation, the optimization path at finite $n$ is governed by the specific gradients $\nabla\mathbf{S}$ induced by the rule.
-The Logarithmic Score, $\mathbf{S}_{Log}(F,y)=-\log f(y)$, exhibits extreme sensitivity to the tails of $\Omega$.
-Since $\nabla\mathbf{S}_{Log}\propto 1/f(y)$, the penalty for low-density observations grows toward infinity.
-At finite $n$, even a single outlier can dominate the gradient, forcing the optimizer to prioritize tail-fitting early, often at the expense of the distribution’s central mass.
-Conversely, the Continuous Ranked Probability Score (CRPS) provides a more robust optimization landscape.
-This is best understood by expressing the CRPS as an integral over pinball losses $\rho_{\tau}$ for all quantiles $\tau\in[0,1]$:
-$\text{CRPS}(F,y)=2\int_{0}^{1}\rho_{\tau}(F^{-1}(\tau),y)d\tau,$ (2)
-where $\rho_{\tau}(u,y)=(y-u)(\tau-\mathbb{I}\{y<u\})$.
-Because the gradient of the pinball loss is bounded (constant in magnitude), the CRPS gradient remains stable regardless of how far $y$ lies in the tails.
-Consequently, at finite $n$, CRPS-based estimation focuses on the global probability mass and is less susceptible to being pulled away by sparse observations in $\Omega$.
-While both rules are asymptotically equivalent in their target, their finite-sample "optimum" differs and also their behavior for tail-sensitivity (LogS) and robust mass-estimation (CRPS).
-1.1 Important Quotes
-  • "It is difficult to understand why statisticians commonly limit their inquiries to averages, and do not revel in more comprehensive views.
-Their souls seem as dull to the charm of variety as that of the native of one of our flat English counties, whose retrospect of Switzerland was that, if its mountains could be thrown into its lakes, two nuisances would be got rid of at once." (Galton, 1889) Kneib et al.
-(2023)
-  • "Despite these well-argued perspectives, there has been little recognition that the common practice of requesting some point forecast, and then evaluating the forecasters by using some (set of) scoring function(s), is not a meaningful endeavor.
-In this article, we develop the perspectives of Murphy and Daan (1985) and Engelberg, Manski, and Williams (2009) and argue that effective point forecasting depends on guidance or directives, which can be given in one of two complementary ways, namely, by disclosing the scoring function ex ante to the forecaster, or by requesting a specific functional of the forecaster’s predictive distribution, such as the mean or a quantile." Gneiting (2011)
-  • "Although every proper scoring rule can identify the true model or the conditional distribution as its minimum in the population sense, they exhibit different inductive biases or regularization effects in finite samples.
-Furthermore, the computational tractability varies substantially depending on the chosen scoring rule" Waghmare and Ziegel (2025)
-1.2 Contributions
-We greatly appreciate the recent improvement of tabular foundation models to move away from point estimates.
-However, current regression benchmarks for Tabular Machine Learning like TabArena or TALENT are still stuck in evaluating "historically important" point estimates only.
-We
-  • propose the enhancement of regression benchmarks to account for probabilistic regression metrics derived from scoring rules, mainly CRPS by discussing a bimodal toy model.
-  • we provide first evaluations of realTabPFNv2.5 and TabICLv2 in terms of proper scoring rules (CRPS and beta energy score).
-  • we show how finetuning with different scoring rules than the one seen during pretraining in a finite-sample setting can enhance realTabPFNv2.5 performance (CRLS, beta energy score)
-  • we show that in theory the choice of a scoring rule can have a significant impact in terms of the inductive bias of the model (having implications for finetuning and benchmarking)Merkle and Steyvers (2013); Waghmare and Ziegel (2025)
-  • Additionally in the outlook, we start the discussion around a potential conflict between optimality of predictions with regard to user-supplied scoring rule and the goal to not adapt a general-purpose foundation model after pretraining.
+Figure 3: The images illustrate the effects of training with CRPS or log score (cross-entropy loss) with the same training data and architecture but a different loss function (CRPS vs.
+log score) at a finite number of samples (1600 train samples, 400 test samples).
+For further description of the models and the data generating process, see the appendix.
+One main takeaway from the toy model in Figure 3 is that mean value estimates may fall in regions without realizations and may (depending on the utility) be completely useless (think about asking: “What could be the next outcome of a six-sided die?” Even if you had a perfect mean value estimate telling you the mean value 3.5, then this would not even be a valid realization).
+Additionally, the usual construction of a prediction interval around the mean value estimate would cover a lot of empty space without samples (just by construction as an interval), which is inferior to estimating a conditional histogram like shown in the figure.
+Interestingly, for the toy dataset, the cross-entropy loss trained model in Figure 3 differs significantly in terms of alignment with the ground truth data generating process, which indicates that at a finite number of samples:
+  • the convergence speed during model training may depend on the chosen loss function, even if both loss functions derive from proper scoring rules
+  • the sample efficiency of the complete training pipeline may depend on the loss function, even if both loss functions derive from proper scoring rules
+In Table 5 in the appendix we observe that different models have different performance metrics depending on the loss function, even under otherwise identical training procedures.
+This behavior has also been observed on other OpenML datasets [^3] See https://github.com/PriorLabs/TabPFN/pull/689 for the pull request proposing CRPS for fine-tuning TabPFN..
 3 Proposed Enhanced Reporting of Proper Scoring Rules in (Tabular) Regression Benchmarks
@@
 3.1 Log Score
-Viewing discretized regression as a multi-class classification problem suggests the application of Cross-Entropy (CE), which corresponds to the log scoring rule which is local.
+Viewing discretized regression as a multi-class classification problem suggests the application of cross-entropy (CE), which corresponds to the log scoring rule, which is local.
 This choice may be optimal if the final business objective is strictly to maximize the probability of the "correct" bin.
 3.2 Continuous Ranked Probability Score (CRPS)
-However, using cross entropy loss (meaning the log scoring rule) would ignore the distance between the prediction and the ground truth (bins are not even treated as having an ordinal order like in $1<2<3$ but more like completely independent concepts like "panda", "tiger", "fish").
-To inject geometric awareness into the gradient, we proposed to use CRPS for finetuning TabPFN in December 2025444https://github.com/PriorLabs/TabPFN/pull/689 and saw CRPS being used in TabICLv2 training in February 2026Qu et al.
+However, using cross-entropy loss (meaning the log scoring rule) would ignore the distance between the prediction and the ground truth (bins are not even treated as having an ordinal order like in $1<2<3$ but more like completely independent concepts like "panda", "tiger", "fish").
+To inject geometric awareness into the gradient, we proposed to use CRPS for fine-tuning TabPFN in December 2025444https://github.com/PriorLabs/TabPFN/pull/689 and saw CRPS being used in TabICLv2 training in February 2026 Qu et al.
 (2026).
@@
 obtained via cumsum of the softmax output of a neural network over the binned support).
-To perform optimum score estimation the expectation value (over all training samples) scoring rule needs to be taken and minimized.
-However, CRPS weights all quantiles of the output space equally.
-This is likely, prohibitive in applications with asymmetric risk structure over the output space (see later discussion also for weighted CRPSGneiting and Ranjan (2011)).
+To perform optimum score estimation, one minimizes the empirical mean of the scoring rule over all training samples.
+However, CRPS weights all quantiles equally, which can be limiting in applications with asymmetric risk over the output space (see the discussion of weighted CRPS below Gneiting and Ranjan (2011)).
 We also think that at finite sample sizes the different scoring rules may differ (under otherwise constant conditions) in convergence speed and in sample efficiency (compare the results of the toy example in Figure 3).
+3.3 Weighted CRPS (wCRPS)
+Accounting for asymmetric risk structures, a modeler can choose to apply the weighted CRPS Gneiting and Ranjan (2011), a family of strictly proper scoring rules that enables emphasizing prediction quality for the left or right tail of a distribution.
+This makes sense when dealing with e.g.
+energy forecasting or forecasting weather events with asymmetric cost.
+In section 4.2 we discuss the practical implications of the choice of the scoring rule for the inductive bias of the model and for fine-tuning.
+However, here we only define it as a potential alternative to the CRPS for fine-tuning and benchmarking (while not using it since weighting would be use-case dependent):
+$L_{wCRPS}\approx\sum_{i=1}^{N}w(x_{i})(F(x_{i})-\mathbbm{1}_{\{x_{i}\geq y\}})^{2}\Delta x_{i},$ (2)
 3.4 Continuous Ranked Logarithmic Scoring Rule (CRLS)
-CRLS is defined by [^5] https://scoringrules.readthedocs.io/en/latest/theory.html
-$\mathrm{CRLS}(F,y)=-\int_{-\infty}^{\infty}\log|F(x)+\mathbf{1}\{y\leq x\}-1|dx.$ (4)
+CRLS (also called "Exceedance Probability Score", EPS) was introduced by Juutilainen et al. (2012) [^5] https://scoringrules.readthedocs.io/en/latest/theory.html
+$\mathrm{CRLS}(F,y)=-\int_{-\infty}^{\infty}\log|F(x)+\mathbf{1}\{y\leq x\}-1|dx,$ (3)
+where a negation is introduced relative to the original paper so that lower scores correspond to better forecasts (minimization convention).
+"The exceedance probability score differs from CRPS so that it evaluates the difference between the binary event and its predicted probability using the logarithm rather than the negative squared difference." Juutilainen et al.
+(2012) "[CRLS] gives more weight to events [than CRPS] that were predicted to be very unlikely.
+[CRLS] differs significantly also from the log scoring rule: [CRLS] is distance sensitive, whereas [Log score] is not." Juutilainen et al.
+(2012) "CRPS is very insensitive to an error in which predicted deviation is too low and that [Log score] is relatively insensitive to an error in which predicted deviation is too high.
+[CRLS] penalizes more on crude errors in predicted mean than LogS or CRPS." Juutilainen et al.
+(2012)
+3.5 Interval Score
+The Interval Score Gneiting and Raftery (2007) targets the $(1-\alpha)$ central prediction interval and is strictly proper, directly answering “how well is the interval estimated?” for a given significance level $\alpha$:
+$S_{\alpha}(F,y)=(u-l)+\frac{2}{\alpha}(l-y)\mathbb{I}\{y<l\}+\frac{2}{\alpha}(y-u)\mathbb{I}\{y>u\}$ (4)
+where $[l,u]$ are the $\alpha/2$ and $1-\alpha/2$ quantiles of the prediction $F$. We report $\mathrm{IS}_{95}$ (i.e. $\alpha=0.05$) as one of our six evaluation metrics.
+This scoring rule highlights that different scoring rules favor optimization in distinct regions of the PMF; in the extreme case of the interval score, only the two boundary quantiles are penalized.
+General-purpose models may therefore fail to achieve optimality for tasks where the cost of error is non-uniform across $\Omega$.
 3.6 Multivariate Enhancements
 For $y\in\Omega\subseteq\mathbb{R}^{d}$ with $d>1$, univariate scores are insufficient for capturing dependence.
-We employ and suggest to benchmark for the Energy Score ($\mathbf{S}_{\beta}$) and Variogram Score ($\mathbf{S}_{VS,p}$) as also done in Landsgesell (2026).
+We employ and suggest benchmarking the Energy Score ($\mathbf{S}_{\beta}$) and Variogram Score ($\mathbf{S}_{VS,p}$), as also done in Landsgesell (2026).
 The Energy Score for $\beta\in(0,2)$ is defined as:
@@
 where $w_{j,k}$ are weights. While $\mathbf{S}_{\beta}$ ensures strict propriety, $\mathbf{S}_{VS,p}$ provides superior sensitivity to misspecified dependencies between components of $y$.
+4 Choice of Scoring Rules for Decision Making
+Different proper scoring rules penalize different types of errors differently and thus induce different inductive biases Waghmare and Ziegel (2025); Merkle and Steyvers (2013). As Gneiting emphasizes:
+“Despite these well-argued perspectives, there has been little recognition that the common practice of requesting some point forecast, and then evaluating the forecasters by using some (set of) scoring function(s), is not a meaningful endeavor.
+[…] effective point forecasting depends on guidance or directives, which can be given in one of two complementary ways, namely, by disclosing the scoring function ex ante to the forecaster, or by requesting a specific functional of the forecaster’s predictive distribution, such as the mean or a quantile.” Gneiting (2011)
+Several lines of work address how to choose the right rule: Oesterheld Oesterheld and Conitzer (2020) formalizes “Decision Scoring Rules”; Johnstone et al.
+Johnstone et al.
+(2011) derive tailored scoring rules from a utility perspective (their orchard weather-forecast example on page 2 is particularly instructive); and Merkle et al.
+Merkle and Steyvers (2013) show empirically that the ranking of forecasters changes with the scoring rule (the Spearman correlation between Brier-score and log-score rankings in their Table 1 is only 0.15).
+We port this idea to the regression setting using the $\beta$-energy score family and show that the optimal model depends on the choice of $\beta$.
+4.1 Finite-Sample Behavior of Proper Scoring Rules
+A proper scoring rule $\mathbf{S}:\mathcal{F}\times\Omega\rightarrow\mathbb{R}$ guarantees that the expected score is minimized by the true distribution $Q\in\mathcal{F}$:
+$\mathbb{E}_{Y\sim Q}[\mathbf{S}(Q,Y)]\leq\mathbb{E}_{Y\sim Q}[\mathbf{S}(F,Y)]\quad\forall F\in\mathcal{F}.$ (7)
+In practice, we minimize the sample average $\widehat{E[\mathbf{S}]}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{S}(F,y_{i})$, whose optimization path at finite $n$ depends on the gradients $\nabla\mathbf{S}$.
+Two properties distinguish the rules used in this paper:
+  • The Logarithmic Score, $\mathbf{S}_{Log}(F,y)=-\log f(y)$, has gradient $\nabla\mathbf{S}_{Log}\propto 1/f(y)$, which is unbounded as $f(y)\to 0$.
+  • The CRPS, expressed as $\text{CRPS}(F,y)=2\int_{0}^{1}\rho_{\tau}(F^{-1}(\tau),y)\,d\tau$ with pinball loss $\rho_{\tau}(u,y)=(y-u)(\tau-\mathbb{I}\{y<u\})$, has a gradient that is bounded in magnitude.
+The unbounded gradient of the log score means that low-density observations contribute disproportionately large updates, whereas CRPS weights all quantile levels equally.
+Both rules are strictly proper and thus share the same population-level minimizer, but their finite-sample optimization dynamics differ due to these gradient properties.
+4.2 Investigating the effect of the choice of scoring rule on the optimal forecast
+For simplicity [^6] While the following argument describes concentrated probabilistic forecasts, the argument remains untouched even if we move to not so extremely concentrated forecasts.
+of the argument, we consider a probabilistic forecast of the following kind:
+$\hat{p}_{i}(Y=y|X=x)=\delta(y-f_{i}(x)),$ (8)
+where $i\in\{A,B,C,D,E\}$ indexes different forecasting models (e.g.
+from different checkpoints of a training) and $\delta()$ is the Dirac delta distribution (which collapses the probability distribution to a "pointy" probability density).
+In this toy model (again for simplicity of the argument) we impose the ground truth
+$p_{g}(Y=y|X=x)=\delta(y-g(x)).$ (9)
+We choose $x\in[-3,3]$777x=np.linspace(-3,3,1000).
+  • $g(x)=\sin(x)$
+  • Model A: Small consistent bias, $f_{A}(x)=g(x)+0.1$
+  • Model B: Perfect with a large localized deviation, $f_{B}(x)=g(x)$ with an additional $2.0$ added to the middle 50 points
+  • Model C: Scaled version of the true function, $f_{C}(x)=0.8\cdot g(x)$
+  • Model D: Shifted version of the true function, $f_{D}(x)=\sin(x+0.2)$
+  • Model E: Simple baseline at zero, $f_{E}(x)=0$
+A practitioner would naturally ask which model is best. We examine the models in Figure 4 for intuition.
+Figure 4: Toy Models for investigating the effect of the choice of scoring rule on the optimal forecast
+The distinct error profiles of the models are visible in Figure 5.
+Some models make small errors everywhere (model A), some models make large errors in a small region (model B), some models make medium-sized errors everywhere (model C) and some models make small errors everywhere but are shifted (model D).
+Figure 5: Absolute errors for the toy models
+Now we can apply different scoring rules (from the $\beta$-energy family) to see which model is optimal for which scoring rule.
+For a fixed choice of $\beta$ (one point on the x-axis in Figure 6) the better model has the lower beta-energy score.
+Figure 6: Decision based on scoring rules for the toy models
+We observe that Model B ranks best for $\beta\lesssim 1$, but for $\beta\gtrsim 1$ the ranking changes, with other models also switching rank (Figure 7).
+Figure 7: Ranking of toy models based on scoring rules
+Fundamentally this behavior arises from different scoring rules rating the severity of errors in Figure 5 differently (and there are even scoring rules which penalize errors depending on the direction of the error, see for example asymmetric CRPS Iacopini et al.
+(2023) which may be important e.g.
+in finance where errors have different severity depending on the direction of the error: making 100 percent more money is less severe than losing 100 percent of money).
+For simple probabilistic point masses like shown here we can even make an analytical statement: The beta-energy score with $\beta=1$ favors minimization of the mean absolute error (i.e.
+prefers estimating a median estimate) while for $\beta=2$ the scoring rule prefers a minimization of mean squared error.
+In Table 1 we also note explicitly the ranking of models changes for three selected $\beta$ values:
+Table 1: Model Rankings for Different $\beta$ Values
+Model Rank ($\beta=0.2$) Rank ($\beta\approx 1.0$) Rank ($\beta=2.0$)
+Model A (Bias) 2 1 1
+Model B (Outlier) 1 4 4
+Model C (Scale) 4 3 3
+Model D (Shift) 3 2 2
+Model E (Zero) 5 5 5
+Derivation.
+The $\beta$-energy score for a forecast distribution $P$ and observation $y\in\mathbb{R}$ is defined as
+$\mathrm{ES}_{\beta}(P,\,y)=\mathbb{E}_{X\sim P}\bigl[|X-y|^{\beta}\bigr]-\tfrac{1}{2}\,\mathbb{E}_{X,\,X^{\prime}\sim P}\bigl[|X-X^{\prime}|^{\beta}\bigr],\qquad 0<\beta\leq 2,$ (10)
+where $X$ and $X^{\prime}$ are independent draws from $P$.
+Now suppose the forecast is a probabilistic point mass, i.e.
+$P=\delta_{\hat{m}}$ for some point estimate $\hat{m}\in\mathbb{R}$.
+Both $X$ and $X^{\prime}$ equal $\hat{m}$ almost surely, so the second term in (10) vanishes and the score reduces to
+$\mathrm{ES}_{\beta}(\delta_{\hat{m}},\,y)=|\hat{m}-y|^{\beta}.$ (11)
+The optimal point estimate $\hat{m}^{*}$ minimises the expected score over observations $y\sim Q$ for some true data distribution $Q$:
+$\hat{m}^{*}=\operatorname*{arg\,min}_{\hat{m}}\;\mathbb{E}_{y\sim Q}\!\left[|\hat{m}-y|^{\beta}\right].$ (12)
+Case $\beta=1$.  Equation (12) becomes $\min_{\hat{m}}\,\mathbb{E}[|\hat{m}-y|]$, i.e. minimisation of the mean absolute error (MAE). The unique minimiser is the median of $Q$, since
+$\frac{\partial}{\partial\hat{m}}\,\mathbb{E}[|\hat{m}-y|]=\Pr(y<\hat{m})-\Pr(y>\hat{m})=0\;\Longleftrightarrow\;\hat{m}=\mathrm{median}(Q).$ (13)
+Case $\beta=2$.  Equation (12) becomes $\min_{\hat{m}}\,\mathbb{E}[(\hat{m}-y)^{2}]$, i.e. minimisation of the mean squared error (MSE). Setting the derivative to zero gives
+$\frac{\partial}{\partial\hat{m}}\,\mathbb{E}\!\left[(\hat{m}-y)^{2}\right]=2\,(\hat{m}-\mathbb{E}[y])=0\;\Longleftrightarrow\;\hat{m}^{*}=\mathbb{E}[y]=\mathrm{mean}(Q).$ (14)
+Hence the $\beta$-energy score with $\beta=2$ is minimised by reporting the mean of the predictive distribution.
+Summary.
+The choice of scoring rule changes the ranking of forecasting models: in this toy model the result follows analytically from the $\beta$-energy score reducing to an $L^{\beta}$ norm.
+Whether this ranking sensitivity carries over to trained distributional models in realistic tabular settings remains an open empirical question, but both the theoretical argument and the empirical literature Waghmare and Ziegel (2025); Merkle and Steyvers (2013) support the view that the choice of scoring rule for pretraining or fine-tuning a tabular foundation model is non-trivial and may affect which use-cases the model is best suited for Oesterheld and Conitzer (2020); Johnstone et al.
+(2011).
+As Waghmare et al.
+summarize:
+“Although every proper scoring rule can identify the true model or the conditional distribution as its minimum in the population sense, they exhibit different inductive biases or regularization effects in finite samples.
+Furthermore, the computational tractability varies substantially depending on the chosen scoring rule.” Waghmare and Ziegel (2025)
+How to best choose a scoring rule for general-purpose tabular foundation model pretraining and how to enable adaptation to business cases with asymmetric error severity remains an open question; we discuss possible routes in section 6.
 5 Results
-We adapt the finetuning code from github to start from the realTabPFNv2.5 checkpoint and implement custom loss functions (Beta energy score (with CRPS as special case), CRLS, and others).
-We fit models on a diverse set of openML datasets for regression as shown in the various result tables.
-For each dataset we limit the amount of samples to randomly selected 3000 samples (random subsampling).
-We then perform 5 fold cross-validation on each subsampled dataset so that we can report performance on holdout data.
-On each split we evaluate a baseline model (always the realTabPFNv2.5) against another model (either a finetuned model or TabICLv2), so that we get paired performance metrics (per split).
-We report several performance metrics ($MAE$, RMSERMSE, $R^{2}$, $CRPS$, CRLSCRLS, $IntervalScore@95$).
-All metrics except R2 are better when they are lower.
-The improvement in $R^{2}$ is reported as absolute difference in $\Delta R^{2}*100$pp (to avoid statistical problems with reporting percent changes on fractions).
-The other improvements are reported as relative change from the baseline.
-For simple interpretation we adopt the sign convention that a positive change is an improvement in the respective metric.
-For the metrics we first compute the change in the metric (at same split) and then compute the average afterwards.
-Table 1: Finetuned TabPFN Regressor Hyperparameters
+We adapt the TabPFN fine-tuning code from GitHub, starting from the realTabPFNv2.5 checkpoint, and implement custom loss functions (the $\beta$-energy score, with CRPS as the $\beta=1$ special case, and CRLS).
+We fit models on a diverse set of OpenML datasets for regression:
+  • Abalone
+  • Ailerons
+  • Ames Housing
+  • Bike Sharing
+  • Black Friday
+  • Brazilian Houses
+  • California Housing
+  • Cpu Small
+  • Diamonds
+  • Digits
+  • Elevators
+  • House 16H
+  • House Sales
+  • Kin8nm
+  • Mercedes Benz
+  • NYC Taxi
+  • Protein
+  • Pol
+  • Puma8NH
+  • Sulfur
+Each dataset is randomly subsampled to at most 3000 instances.
+We then apply 5-fold cross-validation on each subsampled dataset, yielding holdout performance estimates on all available data.
+On each fold, the baseline model (realTabPFNv2.5) and the comparison model (either a fine-tuned variant or TabICLv2) are evaluated on the same held-out split, producing paired performance metrics.
+We report six performance metrics: $\mathrm{MAE}$, $\mathrm{RMSE}$, $R^{2}$, $\mathrm{CRPS}$, $\mathrm{CRLS}$, and $\mathrm{IS}_{95}$ (Interval Score at the 95% level).
+All metrics are lower-is-better, with the exception of $R^{2}$.
+Improvements in $R^{2}$ are expressed as the absolute difference $\Delta R^{2}\times 100$ percentage points (pp), avoiding the statistical ambiguity of relative changes on a fraction.
+Improvements in all other metrics are reported as relative changes with respect to the baseline.
+Throughout, we adopt the sign convention that a positive value denotes an improvement over the baseline, regardless of the metric’s direction.
+For aggregated results, per-fold improvements are computed first and then averaged across folds and datasets.
+The hyperparameters used for fine-tuning the TabPFN regressor are shown in Table 2.
+Table 2: Fine-Tuned TabPFN Regressor Hyperparameters.
+The built-in loss weights (crps_loss_weight etc.) are explicitly set to 0.0 to override the model’s default loss configuration; all weight is instead assigned to the custom scoring-rule implementation.
 Category Hyperparameter Value
@@
 Model Logic average_before_softmax True
-3.1 Beta(1.8)-Energy Score Finetuning realTabPFNv2.5
-We optimize with $\beta=1.8$ energy score and apply early stopping on MAE (changing the metric for early stopping showed strongly varying performance).
-We find improved performance in terms of $R^{2}$:
-Table 2: Comparative Performance and Uncertainty Metrics (Finetuned vs. Base)
-Metric Improvement (mean $\pm$ std) Median W/L/T
-MAE +4.28% $\pm$ 10.06% +1.46% 61/21/8
-RMSE +2.16% $\pm$ 4.39% +0.80% 61/24/5
-$R^{2}$ +1.47pp $\pm$ 3.73pp +0.25pp 52/19/19
-CRPS +2.76% $\pm$ 6.38% +0.79% 61/23/6
-IntervalScore95 +2.24% $\pm$ 13.86% +0.94% 55/33/2
-Table 3: Improvement by Dataset: Finetuned vs. Base
-Dataset MAE Imp. RMSE Imp. $R^{2}$ Imp. CRPS Imp. IS95 Imp.
-Abalone -0.03% +0.47% +0.39pp +0.78% +0.60%
-Ailerons -0.04% -0.24% -0.06pp +0.18% -0.30%
-Bike Sharing +1.49% +1.11% +0.13pp +1.08% -0.49%
-Black Friday +0.16% +0.30% +0.30pp +0.25% -0.74%
-Brazilian Houses -7.33% +3.43% +6.28pp +0.99% +5.64%
-California Housing +0.91% +1.10% +0.48pp +1.25% +1.37%
-Cpu Small +1.38% -0.12% -0.00pp -0.03% +0.92%
-Diamonds +0.61% +1.26% +0.05pp +0.53% -1.46%
-Digits +34.54% +6.29% +0.01pp +8.27% +13.74%
-Elevators +0.02% +0.20% +0.03pp +0.21% +0.82%
-House 16H -0.39% -0.06% -0.08pp -0.47% -0.69%
-Kin8nm +1.83% +2.05% +0.30pp +2.01% +5.18%
-Mercedes Benz +15.71% +6.94% +7.33pp +13.58% +7.27%
-NYC Taxi +8.64% +4.73% +6.68pp +8.18% +3.50%
-Protein +1.44% +0.26% +0.25pp +0.51% +1.12%
-Pol +8.31% +3.30% +3.71pp +5.75% +9.80%
-Puma8NH +7.87% +7.95% +0.77pp +6.85% +7.06%
-Sulfur +1.85% -0.04% -0.02pp -0.17% -13.03%
-Figure 4: Impact of beta=1.8 energy score finetuning on realTabPFNv2.5
-3.2 CRLS-Finetuning realTabPFNv2.5
-We implement the CRLS as a loss in tabpfn and use it here also as evaluation metric for early stopping.
-Table 4: Overall Performance and Uncertainty Improvement (Finetuned vs. Base)
-Metric Improvement (mean $\pm$ std) Median W/L/T
-MAE +0.65% $\pm$ 10.13% +0.21% 54/41/5
-RMSE +1.67% $\pm$ 4.60% +0.96% 73/25/2
-$R^{2}$ +0.94pp $\pm$ 2.40pp +0.33pp 64/20/16
-CRLS +2.27% $\pm$ 7.00% +1.47% 79/18/3
-IntervalScore95 +3.87% $\pm$ 13.15% +1.76% 75/22/3
-Table 5: Detailed Dataset Improvements: Finetuned vs. Base
-Dataset MAE Imp. RMSE Imp. $R^{2}$ Imp. CRLS Imp. IS95 Imp.
-Abalone -0.76% +0.42% +0.35pp +1.52% +2.58%
-Ailerons +1.01% +0.65% +0.17pp +1.52% +0.86%
-Ames Housing -0.21% +1.41% -0.23pp -0.75% -7.59%
-Bike Sharing +1.86% +1.88% +0.21pp +1.19% +1.59%
-Black Friday +0.23% +0.61% +0.61pp +1.09% +0.67%
-Brazilian Houses -24.19% +3.67% +6.72pp -2.45% +23.84%
-California Housing +0.92% +1.94% +0.74pp +2.48% +3.10%
-Cpu Small +0.69% +0.05% +0.00pp +1.14% +1.45%
-Diamonds +0.60% +1.27% +0.05pp +1.01% +1.15%
-Digits +28.23% +7.48% -0.02pp +10.91% +3.09%
-Elevators -0.54% +0.13% +0.02pp +0.34% +0.67%
-House 16H -1.88% +0.42% +0.27pp +0.20% +0.61%
-House Sales -0.71% -1.35% -0.34pp +0.22% -1.02%
-Kin8nm -0.39% -0.77% -0.12pp -0.32% +0.97%
-Mercedes Benz +3.46% +1.14% +1.22pp +2.12% +1.31%
-NYC Taxi +4.71% +3.04% +3.90pp +4.19% +2.08%
-Protein -1.47% +0.51% +0.45pp +1.43% +3.34%
-Pol +7.37% +3.23% +3.64pp +8.15% +13.82%
-Puma8NH +7.24% +7.41% +0.71pp +6.26% +6.37%
-Sulfur -13.24% +0.36% +0.48pp +5.14% +18.50%
-Figure 5: Impact of CRLS finetuning realTabPFNv2.5
-3.3 CRLS performances of TabICLv2 versus realTabPFNv2.5
-Table 6: Aggregate Performance and Uncertainty Improvement (TabICLv2 vs. realTabPFNv2.5)
-Metric Improvement (mean $\pm$ std) Median W/L/T
-MAE +1.41% $\pm$ 12.65% +0.16% 56/48/1
-RMSE +1.85% $\pm$ 10.02% +0.96% 65/34/6
-$R^{2}$ +3.22pp $\pm$ 17.03pp +0.17pp 56/27/22
-CRLS +6.01% $\pm$ 10.95% +3.49% 84/19/2
-IntervalScore95 +5.14% $\pm$ 14.18% +2.15% 70/35/0
-Table 7: Dataset-Specific Improvement Metrics
-Dataset MAE Imp. RMSE Imp. $R^{2}$ Imp. CRLS Imp. IS95 Imp.
-Abalone +0.22% +0.92% +0.76pp +3.37% +4.94%
-Ailerons -1.69% -1.01% -0.29pp +1.55% +0.23%
-Ames Housing +0.37% +8.13% +32.38pp +4.95% -13.51%
-Bike Sharing +0.21% -0.62% -0.06pp +2.83% -0.13%
-Black Friday +0.59% +1.17% +1.18pp +0.34% -2.20%
-Brazilian Houses -13.82% +3.41% +6.13pp +25.02% +31.86%
-California Housing +4.87% +5.25% +1.91pp +8.33% +6.34%
-Cpu Small +1.49% +0.08% +0.00pp +6.17% +3.52%
-Diamonds +1.15% +1.62% +0.06pp +4.33% +4.93%
-Digits +42.92% +21.15% +0.89pp +34.26% +43.52%
-Elevators -0.20% +0.57% +0.10pp +1.55% -0.55%
-House 16H +0.76% +2.92% +1.92pp +5.30% +5.42%
-House Sales -5.00% -10.86% -2.37pp +0.11% -3.01%
-Kin8nm +0.59% -0.29% -0.04pp +1.52% +3.26%
-Mercedes Benz +18.83% +9.57% +9.92pp +20.31% +13.27%
-NYC Taxi -7.26% +3.99% +9.25pp -5.47% -6.48%
-Protein +2.63% +1.71% +1.55pp +4.28% +3.15%
-Pol +10.39% +4.05% +4.56pp +11.22% +7.74%
-Puma8NH -13.29% -13.86% -1.50pp -13.66% -10.89%
-Sulfur -10.13% +1.27% +1.61pp +6.38% +16.39%
-Visualizing Soil -4.08% -0.36% -0.32pp +3.50% +0.08%
-Figure 6: Impact of changing from RealTabPFNv2.5 to TabICLv2: Even stronger outperformance of TabICLv2 over realTabPFNv2.5 on CRLS than seen before by finetuning realTabPFNv2.5 on CRLS
-3.4 CRPS performances of realTabPFNv2.5 and TabICLv2
-Figure 7: Performance Improvement of TabICLv2 over realTabPFNv2.5 on CRPS (and other metrics)
-Table 8: Overall Model Improvement (TabICLv2 vs. realTabPFNv2.5)
-Metric Improvement (mean $\pm$ std) Median W/L/T
-MAE +1.41% $\pm$ 12.65% +0.16% 56/48/1
-RMSE +1.85% $\pm$ 10.02% +0.96% 65/34/6
-$R^{2}$ +3.22pp $\pm$ 17.03pp +0.17pp 56/27/22
-CRPS +2.01% $\pm$ 10.17% +0.73% 59/45/1
-IntervalScore95 +5.14% $\pm$ 14.18% +2.15% 70/35/0
-Table 9: Dataset-Specific Performance Gains
-Dataset MAE Imp. RMSE Imp. $R^{2}$ Imp. CRPS Imp. IS95 Imp.
-Abalone +0.22% +0.92% +0.76pp +0.86% +4.94%
-Ailerons -1.69% -1.01% -0.29pp -2.20% +0.23%
-Ames Housing +0.37% +8.13% +32.38pp -1.81% -13.51%
-Bike Sharing +0.21% -0.62% -0.06pp -0.75% -0.13%
-Black Friday +0.59% +1.17% +1.18pp +0.61% -2.20%
-Brazilian Houses -13.82% +3.41% +6.13pp +3.63% +31.86%
-California Housing +4.87% +5.25% +1.91pp +5.84% +6.34%
-Cpu Small +1.49% +0.08% +0.00pp +1.82% +3.52%
-Diamonds +1.15% +1.62% +0.06pp +1.45% +4.93%
-Digits +42.92% +21.15% +0.89pp +28.19% +43.52%
-Elevators -0.20% +0.57% +0.10pp -0.17% -0.55%
-House 16H +0.76% +2.92% +1.92pp +1.82% +5.42%
-House Sales -5.00% -10.86% -2.37pp -4.79% -3.01%
-Kin8nm +0.59% -0.29% -0.04pp +0.20% +3.26%
-Mercedes Benz +18.83% +9.57% +9.92pp +20.91% +13.27%
-NYC Taxi -7.26% +3.99% +9.25pp -8.96% -6.48%
-Protein +2.63% +1.71% +1.55pp +2.79% +3.15%
-Pol +10.39% +4.05% +4.56pp +6.73% +7.74%
-Puma8NH -13.29% -13.86% -1.50pp -15.15% -10.89%
-Sulfur -10.13% +1.27% +1.61pp +1.78% +16.39%
-Visualizing Soil -4.08% -0.36% -0.32pp -0.56% +0.08%
-4 Choice of Scoring Rules for Decision Making
-The choice of a scoring rule matters because different scoring rules have different inductive biases Waghmare and Ziegel (2025); Merkle and Steyvers (2013).
-This is the reason why the scoring rule used for solving a business problem should be chosen carefully:
-  • Oesterheld and Conitzer (2020) describes "Decision Scoring Rules"
-  • Winkler et al. describe tailored scoring rules for decision making from a utility perspectiveJohnstone et al. (2011).
-  • The very insightful introduction example in Johnstone et al.
-(2011), page 2, is about making "useful" probabilistic weather forecasts for a specific use-case (an orchard) and we drew a lot of inspiration reading it - this example is very intuitive we suggest to read itJohnstone et al.
-(2011).
-  • Merkle et al.Merkle and Steyvers (2013) in their paper "Choosing a strictly proper scoring rule" describe an important effect: The "optimal forecaster" depends on the choice of a scoring rule (while all scoring rules are individually minimized by a forecast which is equal to the true distribution).
-In other words, different scoring rules penalize different types of errors differently and thus the optimal forecast (the one which minimizes the expected score) can differ based on the scoring rule used.
-Merkle already points this out and finds that different forecasters are ranked very differently by different strictly proper scoring rules.
-They for example find in table 1 that the spearman correlation for the ranking of forecasters between the brier score and the log score is only 0.15.
-They then go on and investigate a parametrizable family of scoring rules and see how the ranking generally depends on the choice of the parameters.
-For this article we follow this idea and port it to proper scoring rules which can be used in the regression setting.
-We choose the $\beta$-energy score family and show that the optimal model depends on the choice of $\beta$.
-To show this and for keeping the arguments as simple as possible we consider
-4.1 Invesigating the effect of the choice of scoring rule on the optimal forecast
-For simplicity [^6] While the following argument describes concentrated probabilistic forecasts, the argument remains untouched even if we move to not so extremely concentrated forecasts.
-of the argument, we consider a probabilistic forecast of the following kind:
-$\hat{p}_{i}(Y=y|X=x)=\delta(y-f_{i}(x)),$ (7)
-where $i\in\{A,B,C,D,E\}$ indexes different forecasting models (e.g.
-from different checkpoints of a training) and $\delta()$ is the Dirac delta distribution (which collapses the probability distribution to a "pointy" probability density).
-In this toy model (again for simplicity of the argument) we impose the ground truth
-$p_{g}(Y=y|X=x)=\delta(y-g(x)).$ (8)
-We choose $x\in[-3,3]$ (np.linspace(-3,3,1000)).
-  • $g(x)=\sin(x)$
-  • Model A: Small consistent bias, $f_{A}(x)=g(x)+0.1$
-  • Model B: Perfect with a large localized deviation, $f_{B}(x)=g(x)$ with an additional $2.0$ added to the middle 50 points
-  • Model C: Scaled version of the true function, $f_{C}(x)=0.8\cdot g(x)$
-  • Model D: Shifted version of the true function, $f_{D}(x)=\sin(x+0.2)$
-  • Model E: Simple baseline at zero, $f_{E}(x)=0$
-A typical user would now wonder which model is best? To get an intuition, we look at the models in Figure 8.
-Figure 8: Toy Models for investigating the effect of the choice of scoring rule on the optimal forecast
-Obviously the different models make different erros which can be seen in the absolute error plot in Figure 9.
-Some models make small errors everywhere (model A), some models make large errors in a small region (model B), some models make medium sized errors everywhere (model C) and some models make small errors everywhere but are shifted (model D).
-Figure 9: Absolute errors for the toy models
-Now we can apply different scoring rules (from the $\beta$-energy family) to see which model is optimal for which scoring rule.
-For a fixed choice of $\beta$ (one point on the x-axis in Figrue 10) the better model has the lower beta energy score.
-Figure 10: Decision based on scoring rules for the toy models
-We observe that while the model B is the best ranked model for each $\beta\lesssim 1$ for $1\lesssim\beta$ this changes and also other models switch rank at other values of $\beta$, see figure 11.
-Figure 11: Ranking of toy models based on scoring rules
-Fundamentally this behavior arises from different scoring rules rating the severity of errors in Figure 9 differently (and there is even scoring rules wich penalize errors depending on the direction of the error, see for example asymmetric CRPSIacopini et al.
-(2023) which may be important e.g.
-in finance where errors have different severity depending on the direction of the error: making 100 percent more money is less severe than loosing 100 percent of money).
-For simple probabilistic point masses like shown here we can even make an analytical statement: The beta energy score with $\beta=1$ favors minimization of the mean absolute error (i.e.
-prefers estimating a median estimate) while for $\beta=2$ the scoring rule prefers a minimization of mean squared error.
-Derivation.
-The $\beta$-energy score for a forecast distribution $P$ and observation $y\in\mathbb{R}$ is defined as
-$\mathrm{ES}_{\beta}(P,\,y)=\mathbb{E}_{X\sim P}\bigl[|X-y|^{\beta}\bigr]-\tfrac{1}{2}\,\mathbb{E}_{X,\,X^{\prime}\sim P}\bigl[|X-X^{\prime}|^{\beta}\bigr],\qquad 0<\beta\leq 2,$ (9)
-where $X$ and $X^{\prime}$ are independent draws from $P$.
-Now suppose the forecast is a probabilistic point mass, i.e.
-$P=\delta_{\hat{m}}$ for some point estimate $\hat{m}\in\mathbb{R}$.
-Both $X$ and $X^{\prime}$ equal $\hat{m}$ almost surely, so the second term in (9) vanishes and the score reduces to
-$\mathrm{ES}_{\beta}(\delta_{\hat{m}},\,y)=|\hat{m}-y|^{\beta}.$ (10)
-The optimal point estimate $\hat{m}^{*}$ minimises the expected score over observations $y\sim Q$ for some true data distribution $Q$:
-$\hat{m}^{*}=\operatorname*{arg\,min}_{\hat{m}}\;\mathbb{E}_{y\sim Q}\!\left[|\hat{m}-y|^{\beta}\right].$ (11)
-Case $\beta=1$.  Equation (11) becomes $\min_{\hat{m}}\,\mathbb{E}[|\hat{m}-y|]$, i.e. minimisation of the mean absolute error (MAE). The unique minimiser is the median of $Q$, since
-$\frac{\partial}{\partial\hat{m}}\,\mathbb{E}[|\hat{m}-y|]=\Pr(y<\hat{m})-\Pr(y>\hat{m})=0\;\Longleftrightarrow\;\hat{m}=\mathrm{median}(Q).$ (12)
-Case $\beta=2$.  Equation (11) becomes $\min_{\hat{m}}\,\mathbb{E}[(\hat{m}-y)^{2}]$, i.e. minimisation of the mean squared error (MSE). Setting the derivative to zero gives
-$\frac{\partial}{\partial\hat{m}}\,\mathbb{E}\!\left[(\hat{m}-y)^{2}\right]=2\,(\hat{m}-\mathbb{E}[y])=0\;\Longleftrightarrow\;\hat{m}^{*}=\mathbb{E}[y]=\mathrm{mean}(Q).$ (13)
-Hence the $\beta$-energy score with $\beta=2$ is minimised by reporting the mean of the predictive distribution.
-The choice of the scoring rule determines how different types of errors are penalized and thus the optimal forecast (the one which minimizes the expected score) can differ based on the scoring rule used.
-This is true, despite every scoring rule being strictly proper (i.e.
-every scoring rule is minimized by the true distribution).
-The reason is that different scoring rules penalize different types of errors differently and thus the optimal forecast (the one which minimizes the expected score) can differ based on the scoring rule used.
-We leave it as an open question how to best choose a scoring rule for general purpose tabular foundation model pretraining and how to enable adaptation to arbitrary business case with high severity for certain errors (like in finance or pharmaceutical production).
-We describe some possible routes to explore in the outlook section 7.
-5 Discussion
-Across the datasets evaluated, finetuning with beta-energy score ($\beta=1.8$) and CRLS both tend to improve realTabPFNv2.5 on the majority of metrics and datasets (as indicated by the W/L/T counts in Tables 2 and 4).
-However, improvements are modest in median terms (typically $2\%$) and vary considerably across datasets.
-The aggregate results further suggest that TabICLv2 may be a stronger probabilistic model than realTabPFNv2.5 in terms of CRPS, CRLS, and Interval Score, with TabICLv2 winning on the majority of evaluated datasets (Tables 6, 8).
+We fine-tune TabPFN with multiple scoring rules; here we report results for CRLS and the $\beta=1.8$ energy score, evaluating each variant against the base model (realTabPFNv2.5).
+  • We optimize with $\beta=1.8$ energy score and apply early stopping on MAE (changing the metric for early stopping showed strongly varying performance, because the early stopping metric is also involved in model selection through picking specific checkpoints).
+  • We implement CRLS as a loss function in TabPFN and use it also as the early-stopping metric.
+  • We additionally report TabICLv2 results against the base model to contextualize the fine-tuned variants relative to another strong baseline.
+Aggregated paired performance improvements above the baseline are reported in Table 3.
+Table 3: Aggregated paired performance improvement across all model comparisons.
+All improvements are relative to the realTabPFNv2.5 baseline.
+Dashes indicate metrics not evaluated for that comparison.
+Bold values indicate the best performer for that specific column/metric combination.
+W/L/T counts the number of datasets where the model exceeded the baseline by more than a small $\varepsilon$ (Win), the baseline was better by more than $\varepsilon$ (Loss), or the absolute difference was at most $\varepsilon$ (Tie).
+Beta(1.8)-Energy FT CRLS FT TabICLv2
+Metric mean $\pm$ std Median W/L/T mean $\pm$ std Median W/L/T mean $\pm$ std Median W/L/T
+$\mathrm{MAE}$ +4.28% $\pm$ 10.06% +1.46% 61/21/8 +0.65% $\pm$ 10.13% +0.21% 54/41/5 +1.41% $\pm$ 12.65% +0.16% 56/48/1
+$\mathrm{RMSE}$ +2.16% $\pm$ 4.39% +0.80% 61/24/5 +1.67% $\pm$ 4.60% +0.96% 73/25/2 +1.85% $\pm$ 10.02% +0.96% 65/34/6
+$R^{2}$ +1.47pp $\pm$ 3.73pp +0.25pp 52/19/19 +0.94pp $\pm$ 2.40pp +0.33pp 64/20/16 +3.22pp $\pm$ 17.03pp +0.17pp 56/27/22
+$\mathrm{CRPS}$ +2.76% $\pm$ 6.38% +0.79% 61/23/6 +2.01% $\pm$ 10.17% +0.73% 59/45/1
+$\mathrm{CRLS}$ +2.27% $\pm$ 7.00% +1.47% 79/18/3 +6.01% $\pm$ 10.95% +3.49% 84/19/2
+$\mathrm{IS}_{95}$ +2.24% $\pm$ 13.86% +0.94% 55/33/2 +3.87% $\pm$ 13.15% +1.76% 75/22/3 +5.14% $\pm$ 14.18% +2.15% 70/35/0
+The results on the dataset level are reported in the appendix 8.2.
+The empirical results demonstrate a clear performance trade-off between point-estimate accuracy and probabilistic calibration:
+Point Estimation Superiority:
+The Beta(1.8)-Energy FT variant is the strongest performer for absolute error reduction.
+It achieves the highest mean improvement in $\mathrm{MAE}$ ($4.28\%$) and $\mathrm{CRPS}$ ($2.76\%$), suggesting it is the most effective model when the primary objective is minimizing the distance between predictions and ground truth.
+Consistency and Robustness:
+CRLS FT exhibits high consistency across diverse datasets.
+While its mean improvements are more conservative, it secures the highest number of wins in $\mathrm{RMSE}$ ($73/25/2$) and $R^{2}$ ($64/20/16$).
+This indicates that the CRLS fine-tuning objective successfully mitigates the impact of outliers, offering a more stable "frequentist" profile.
+Probabilistic Calibration:
+TabICLv2 dominates the interval-based and probabilistic scoring metrics.
+It provides the most significant gains in $\mathrm{CRLS}$ ($6.01\%$) and $\mathrm{IS}_{95}$ ($5.14\%$).
+Furthermore, its superior mean $R^{2}$ improvement ($3.22\text{pp}$) suggests a high capacity for capturing complex, non-linear dependencies while maintaining well-calibrated uncertainty quantification.
+We see that on average Beta(1.8)-Energy FT is beneficial for tasks requiring the most accurate point predictions in terms of MAE, while TabICLv2 is the superior choice for applications where high-quality confidence intervals and density estimation are critical or when a high mean $R^{2}$ is to be achieved.
+However, when median $R^{2}$ performance is to be optimized, fine-tuning with more tail-sensitive metrics like CRLS is beneficial.
+Further differences between TabICLv2 and realTabPFNv2.5 may be due to other factors like
+  • different prior data generating processes
+  • different architecture with different memory scaling inducing different inductive biases
+We note that frequent differences in $\mathrm{RMSE}$ and $R^{2}$ performance are due to relative improvement aggregation and different noise levels across datasets.
+Since $R^{2}$ is a scale-independent metric (effectively a per-dataset scaled $\mathrm{RMSE}$), it provides a more balanced view of performance across tasks with varying target variances, whereas mean improvements in $\mathrm{RMSE}$ can remain sensitive to the specific error scales of the underlying datasets.
+5.1 TabICLv2 vs RealTabPFNv2.5 Single Dataset Results
+The per-dataset results are shown in the appendix.
+The performance difference between TabICLv2 and realTabPFNv2.5 is heavily concentrated in specific datasets, revealing the strengths and edge cases of each PFN:
+Point Estimation (MAE & RMSE):
+The most substantial gains are observed in the Digits ($42.92\%$ MAE) and Mercedes Benz ($18.83\%$ MAE) datasets, where TabICLv2 appears to capture the underlying structure significantly better than the baseline.
+Conversely, TabICLv2 struggles with Puma8NH ($13.29\%$ MAE) and Brazilian Houses ($13.82\%$ MAE).
+Interestingly, in Brazilian Houses, the model loses in MAE but wins in RMSE ($3.41\%$), suggesting that while its median predictions may be further off, it is better at avoiding the large squared errors that penalize RMSE.
+Goodness-of-Fit ($R^{2}$):
+The improvement in $R^{2}$ is dominated by Ames Housing ($32.38\text{pp}$), which is a massive outlier compared to the rest of the benchmark.
+Significant positive shifts also occur in Mercedes Benz ($9.92\text{pp}$) and NYC Taxi ($9.25\text{pp}$).
+These high-impact gains are contrasted by minor losses in the House Sales ($2.37\text{pp}$) and Puma8NH ($1.50\text{pp}$) datasets, indicating that TabICLv2’s inductive bias is particularly potent for high-dimensional feature spaces like housing and automotive pricing.
+Likelihood Calibration (CRLS):
+CRLS improvements are nearly universal, but the magnitude is extreme in Digits ($34.26\%$) and Brazilian Houses ($25.02\%$).
+Even in datasets where point metrics are neutral or slightly negative, such as Cpu Small and Sulfur, the CRLS improves by $6.17\%$ and $6.38\%$ respectively.
+This demonstrates that even if the prediction centers are similar, TabICLv2 is far superior at modeling the shape of the uncertainty.
+Global Error (CRPS):
+The CRPS metric reinforces the dominance of TabICLv2 on Digits ($28.19\%$) and Mercedes Benz ($20.91\%$).
+Notably, in NYC Taxi, the model suffers a $8.96\%$ drop in CRPS, which is its second-largest loss after Puma8NH ($15.15\%$).
+This suggests that for certain temporal or high-frequency spatial datasets, the realTabPFNv2.5 prior remains highly competitive.
+Interval Estimation (IS95):
+The Brazilian Houses dataset shows the most dramatic improvement here ($31.86\%$), alongside Digits ($43.52\%$) and Sulfur ($16.39\%$).
+The only major degradation occurs in Ames Housing ($13.51\%$) and Puma8NH ($10.89\%$).
+The Ames result is particularly curious: while the model provides the best $R^{2}$ (explained variance) for this dataset, it provides poor interval coverage, suggesting the model is “overconfident” about its high-accuracy predictions on this specific task.
+This illustrates a general pitfall: a model can sharpen its point predictions (improving $R^{2}$) while simultaneously narrowing its predictive intervals too aggressively, producing intervals that are too tight and thus under-cover the true outcome.
+It is precisely this kind of trade-off that motivates reporting both point-estimate metrics and proper scoring rules such as the Interval Score side by side.
+Summary: Overall, fine-tuning realTabPFNv2.5 as well as switching to TabICLv2 has a positive impact on several performance metrics compared to the baseline.
+However, as seen in Table 3, each model exhibits distinct strengths aligned with its training objective or coming from different pretraining datasets and architecture.
+Fine-tuning with beta-energy score ($\beta=1.8$) and CRLS both improve realTabPFNv2.5 on the majority of metrics and datasets (as indicated by the W/L/T counts in Tables 6 and 8), though improvements are modest in median terms (typically $2\%$) and vary considerably across datasets.
+TabICLv2 appears to be a stronger probabilistic model than realTabPFNv2.5 in terms of CRPS, CRLS, and $\mathrm{IS}_{95}$, winning on the majority of evaluated datasets (Tables 10, 12).
 The margin varies substantially across datasets (e.g.
 large gains on Mercedes Benz and Pol, but losses on Puma8NH and House Sales).
-Additional experiments examining how finetuning performance scales with dataset size would be valuable to better understand whether the observed differences are robust across data regimes.
-We illustrate in section 4.1 that the choice of scoring rule can change the ranking of forecasting models — at least in the considered toy setting of Dirac-delta forecasts over a smooth ground truth function.
-Under this simplification, the result follows analytically from the $\beta$-energy score reducing to an $L^{\beta}$ norm.
-Whether this ranking sensitivity carries over to actual trained distributional models in realistic tabular settings remains an open empirical question.
-That said, the theoretical argument and the empirical literature Waghmare and Ziegel (2025); Merkle and Steyvers (2013) both support the view that the choice of scoring rule for pretraining or finetuning a tabular foundation model is non-trivial and may affect which use-cases the model is best suited for Oesterheld and Conitzer (2020); Johnstone et al.
-(2011).
-6 Limitations: Rare or Extreme Events
-We want to highlight that optimal scoring rule estimation (minimizing/maximizing the expected value of a scoring rule) is inherently not well suited for eliciting properties which are only realized in rare events (fall in the area where no data was observed or in 1D are in the tail of a distribution).
-This was well argued in recent years for the case of one dimensional distributions but equally carries over to higher dimensions.
-Lerch et al.
-(2017); Brehmer and Strokorb (2019) Inherently, this problem is related to epistemic uncertainty (uncertainty due to not having training data).
-Based on this insight also PFNs (trained to optimize mean scoring rules) should be applied with care if risk associated mainly with rare (tail) events or events outside the support of the examples in the supplied context.
+We attribute these complementary strength profiles in part to the fundamental inductive biases imposed through different scoring-rule-based loss functions: as discussed in section 4, optimizing a Beta-Energy score with $\beta$ closer to $2$ favors MSE minimization, and training with any one of the infinitely many possible scoring rules dictates how specific approximation errors are penalized.
 6 Outlook
-While we discussed how to properly measure the goodness of a probabilistic forecast - namely by applying proper scoring rules, we highlight, that this discussion is not finished.
-Gneiting pointed out in Gneiting and Ranjan (2011) that situations with asymmetric risk (finance, weather forecasting etc.) may require to optimize for specific point estimates (via scoring functions suitable to elicit a specific point estimate).
-In such a case, we wonder how tabular foundation models would adapt to specific user supplied scoring functions (so that optimality is guaranteed).
-We point out a conflict of
-  • a foundation model which is not altered by a user at all after pretraining is finished
-  • the need to take into account that the optimal scoring function depends on the use-case at hand (must be supplied by the user since the user intention dictates the utility).
-We see that Fine-Tuning of foundation models could be one route to enable optimal predictions for arbitrary scoring functions.
-Another route could be to enable task-token conditional probability prediction in case of asymmetric risk structures (a user would prompt the model with a desired scoring rule which should be optimal, but this behaviour would need to be pretrained).
-Such a task-token conditional probability prediction could output probability estimates which are optimal under some parametrizable family of scoring rules (like e.g.
-the beta energy score family).
-However there is infinitely many scoring rules and different families of scoring rules so that it is unlcear if pre-training with taks-tokens can in general adapt to arbitrary scoring rules as objectives.
-Potentially, tabular foundation models for high stakes applications need to make sensible choices for task conditioning for certain applications (finance, pharma production, etc.).
-7.1 Example: Weighted CRPS (wCRPS)
-Accounting for asymmetric risk structures, a modeler can choose to apply the different versions of weighted CRPSGneiting and Ranjan (2011) - a family of strictly proper scoring rules which enables e.g.
-emphasizing prediction quality for the left/right tail of a distribution which makes sense when dealing with e.g.
-energy forecasting or forecasting weather events with asymmetric cost.
-$L_{wCRPS}\approx\sum_{i=1}^{N}w(x_{i})(F(x_{i})-\mathbbm{1}_{\{x_{i}\geq y\}})^{2}\Delta x_{i},$ (14)
-7.2 Example: Interval Score
-The Interval ScoreGneiting and Raftery (2007) targets the $(1-\alpha)$ central interval and is strictly proper, directly answering "how well is the interval estimated?" for a given significance level $\alpha$:
-$S_{\alpha}(F,y)=(u-l)+\frac{2}{\alpha}(l-y)\mathbb{I}\{y<l\}+\frac{2}{\alpha}(y-u)\mathbb{I}\{y>u\}$ (15)
-where $[l,u]$ are the $\alpha/2$ and $1-\alpha/2$ quantiles of the prediction $F$.
-This specific scoring rule highlights that different scoring rules favor optimization in distinct regions of the probability mass function (PMF); in the extreme case of the interval score, the two boundary quantiles are of primary importance.
-Because we hypothesize that each scoring rule aligns with a specific utility, general-purpose models may fail to achieve optimality for specialized downstream tasks where the cost of error is non-uniform across the domain $\Omega$.
+Proper scoring rules provide a principled framework for evaluation and should be reported in benchmarks on tabular regression.
+However, several open questions remain for tabular foundation models: Gneiting pointed out in Gneiting and Ranjan (2011) that asymmetric risk settings (finance, weather forecasting, etc.) may require optimizing for specific point functionals via tailored scoring rules.
+A fundamental tension arises between:
+  • a general-purpose foundation model that is not modified after pretraining, and
+  • the requirement that the optimal scoring rule be use-case-specific and therefore supplied by the user.
+Two potential routes to resolve this tension are: (i) fine-tuning on a user-supplied scoring rule at deployment time, or (ii) task-token conditioning, where the model is prompted with a desired scoring rule or parameter of some scoring rule family for which it was trained to produce optimal forecasts under that rule.
+The second approach faces a fundamental challenge: there exist infinitely many proper scoring rules Gneiting and Raftery (2007), spanning infinitely many parametric families, so it is unclear whether pretraining with task tokens can generalize to arbitrary scoring objectives.
+High-stakes applications (finance, pharmaceutical production) with asymmetric risk may therefore require model variants pretrained with carefully chosen task-specific conditioning or fine-tuning.
+Finally, a structural limitation of all scoring-rule-based approaches is that they are inherently ill-suited for eliciting properties realized only in rare or tail events Lerch et al.
+(2017); Brehmer and Strokorb (2019).
+Where training data is absent (e.g.
+in tail regions), no scoring rule can compensate; this is fundamentally a problem of epistemic uncertainty.
+PFNs trained to optimize expected scoring rules should therefore be applied with care when risk is concentrated in the tails or outside the support of the supplied context.
 7 Acknowledgement
@@
 8 Appendix
-9.1 Toy-Models to investigate finite sample effects
+8.1 Toy Models to investigate finite sample effects
 8.1.1 Data Generating Process
@@
 Probabilistic Classification Models These models predict probabilities across $N=50$ discrete bins of the $y$-variable.
-The neural networks are optimized using the Adam optimizer and are simple multi-layer perceptron (MLP) featuring an input layer, two hidden layers with 64 neurons each using $\tanh$ activations, and a Softmax output layer for $n$ bins.
+The neural networks are optimized using the Adam optimizer and are simple multi-layer perceptrons (MLPs) featuring an input layer, two hidden layers with 64 neurons each using $\tanh$ activations, and a Softmax output layer for $n$ bins.
 Hyperparameter PyTorch CRPS Model PyTorch Cross-Entropy Model
@@
 Training Parameters
-Loss Function RPS Loss CrossEntropyLoss
+Loss Function CRPS Loss CrossEntropyLoss
 Optimizer Adam Adam
@@
   • PyTorch Cross-Entropy (CE) Model: A feed-forward neural network trained with cross-entropy loss (log score), aiming to maximize the probability of the true bin.
-  • PyTorch Ranked Probability Score (RPS) Model: A neural network trained directly to minimize the expected CRPS, which optimizes the accuracy of the cumulative distribution function (CDF).
-  • XGBoost RPS Model: An XGBoost vectorleaf classifier utilizing a custom objective function that directly optimizes the Continuous Ranked Probability Score
+  • PyTorch CRPS Model: A neural network trained directly to minimize the expected CRPS, which optimizes the accuracy of the cumulative distribution function (CDF).
+  • XGBoost CRPS Model: An XGBoost vector-leaf classifier that uses a custom objective function to directly optimize the Continuous Ranked Probability Score:
     1import numpy as np
@@
     6 """
-    7 Improved RPS Objective for XGBoost.
+    7 Improved CRPS Objective for XGBoost.
     8 Optimized for multi-class flattening and numerical stability.
@@
     45
-    46def rps_metric(preds, dtrain):
-    47 """Evaluation metric for RPS."""
-    48 labels = dtrain.get_label().astype(int)
-    49 num_samples = labels.shape[0]
-    50 num_bins = preds.size // num_samples
-    51
-    52 preds = preds.reshape(num_samples, num_bins)
-    53 probs = F.softmax(torch.tensor(preds), dim=1).numpy()
-    54
-    55 target_one_hot = np.zeros_like(probs)
-    56 target_one_hot[np.arange(num_samples), labels] = 1
-    57
-    58 pred_cdf = np.cumsum(probs, axis=1)
-    59 target_cdf = np.cumsum(target_one_hot, axis=1)
-    60
-    61 rps = np.mean(np.sum((pred_cdf - target_cdf)**2, axis=1))
-    62 return ’RPS’, rps
-    63
-    64import xgboost as xgb
-    65
-    66# 1. Convert training data
-    67dtrain = xgb.DMatrix(X_train, label=y_train_labels)
-    68
-    69# 2. Parameters
-    70params = {
-    71 ’objective’: ’rank:pairwise’,
+    46import xgboost as xgb
+    47
+    48# 1. Convert training data
+    49dtrain = xgb.DMatrix(X_train, label=y_train_labels)
+    50
+    51# 2. Parameters
+    52params = {
+    53 ’objective’: ’rank:pairwise’, #overwritten later by custom objective
     54 ’eval_metric’: ’rmse’,
@@
     64 ’alpha’: 0.5,
-    83}
-    84
-    85# 3. Train model
-    86num_boost_round = 100
-    87xgb_rps_model = xgb.train(
+    65}
+    66
+    67# 3. Train model
+    68num_boost_round = 100
+    69xgb_crps_model = xgb.train(
     70 params,
@@
     73 obj=crps_objective,
-    92)
+    74)
 8.1.3 Model Performance Comparison
-The performance is evaluated using Mean Squared Error (MSE), Mean Absolute Error (MAE), and the Ranked Probability Score (CRPS).
+The performance is evaluated using Mean Squared Error (MSE), Mean Absolute Error (MAE), and the Continuous Ranked Probability Score (CRPS).
 Table 5: Performance Metrics Comparison
@@
   • Regression Failure: Least Squares and Least Absolute Deviations models struggle to capture the underlying distribution, as point estimates often fall in the low-probability "valley" between modes.
-  • RPS Advantage: Both the PyTorch and XGBoost models trained with the custom CRPS objective provide superior probabilistic forecasts.
+  • CRPS Advantage: Both the PyTorch and XGBoost models trained with the custom CRPS objective provide superior probabilistic forecasts.
   • Conclusion: While MSE and MAE (as specific scoring functions, see Gneiting and Ranjan (2011)) provide insights into point prediction, CRPS is the appropriate metric for evaluating complex conditional distributions.
-The XGBoost model with the custom CRPS objective demonstrated the most robust performance.
+8.2 Detailed Results
+8.3 Beta(1.8)-Energy Score Fine-Tuning realTabPFNv2.5
+We find improved performance in terms of $R^{2}$:
+Table 6: Comparative Performance Metrics (Fine-Tuned vs.
+Base).
+W/L/T counts the number of datasets where the fine-tuned model exceeded the baseline by more than a small $\varepsilon$ (Win), the baseline was better by more than $\varepsilon$ (Loss), or the absolute difference was at most $\varepsilon$ (Tie).
+Metric Improvement (mean $\pm$ std) Median W/L/T
+MAE +4.28% $\pm$ 10.06% +1.46% 61/21/8
+RMSE +2.16% $\pm$ 4.39% +0.80% 61/24/5
+$R^{2}$ +1.47pp $\pm$ 3.73pp +0.25pp 52/19/19
+CRPS +2.76% $\pm$ 6.38% +0.79% 61/23/6
+IntervalScore95 +2.24% $\pm$ 13.86% +0.94% 55/33/2
+Table 7: Improvement by Dataset: Fine-Tuned vs. Base
+Dataset MAE Imp. RMSE Imp. $R^{2}$ Imp. CRPS Imp. IS95 Imp.
+Abalone -0.03% +0.47% +0.39pp +0.78% +0.60%
+Ailerons -0.04% -0.24% -0.06pp +0.18% -0.30%
+Bike Sharing +1.49% +1.11% +0.13pp +1.08% -0.49%
+Black Friday +0.16% +0.30% +0.30pp +0.25% -0.74%
+Brazilian Houses -7.33% +3.43% +6.28pp +0.99% +5.64%
+California Housing +0.91% +1.10% +0.48pp +1.25% +1.37%
+Cpu Small +1.38% -0.12% -0.00pp -0.03% +0.92%
+Diamonds +0.61% +1.26% +0.05pp +0.53% -1.46%
+Digits +34.54% +6.29% +0.01pp +8.27% +13.74%
+Elevators +0.02% +0.20% +0.03pp +0.21% +0.82%
+House 16H -0.39% -0.06% -0.08pp -0.47% -0.69%
+Kin8nm +1.83% +2.05% +0.30pp +2.01% +5.18%
+Mercedes Benz +15.71% +6.94% +7.33pp +13.58% +7.27%
+NYC Taxi +8.64% +4.73% +6.68pp +8.18% +3.50%
+Protein +1.44% +0.26% +0.25pp +0.51% +1.12%
+Pol +8.31% +3.30% +3.71pp +5.75% +9.80%
+Puma8NH +7.87% +7.95% +0.77pp +6.85% +7.06%
+Sulfur +1.85% -0.04% -0.02pp -0.17% -13.03%
+Figure 8: Impact of beta=1.8 energy score fine-tuning on realTabPFNv2.5
+8.4 CRLS Fine-Tuning realTabPFNv2.5
+Table 8: Overall Performance Improvement (Fine-Tuned vs. Base)
+Metric Improvement (mean $\pm$ std) Median W/L/T
+MAE +0.65% $\pm$ 10.13% +0.21% 54/41/5
+RMSE +1.67% $\pm$ 4.60% +0.96% 73/25/2
+$R^{2}$ +0.94pp $\pm$ 2.40pp +0.33pp 64/20/16
+CRLS +2.27% $\pm$ 7.00% +1.47% 79/18/3
+IntervalScore95 +3.87% $\pm$ 13.15% +1.76% 75/22/3
+Table 9: Detailed Dataset Improvements: Fine-Tuned vs. Base
+Dataset MAE Imp. RMSE Imp. $R^{2}$ Imp. CRLS Imp. IS95 Imp.
+Abalone -0.76% +0.42% +0.35pp +1.52% +2.58%
+Ailerons +1.01% +0.65% +0.17pp +1.52% +0.86%
+Ames Housing -0.21% +1.41% -0.23pp -0.75% -7.59%
+Bike Sharing +1.86% +1.88% +0.21pp +1.19% +1.59%
+Black Friday +0.23% +0.61% +0.61pp +1.09% +0.67%
+Brazilian Houses -24.19% +3.67% +6.72pp -2.45% +23.84%
+California Housing +0.92% +1.94% +0.74pp +2.48% +3.10%
+Cpu Small +0.69% +0.05% +0.00pp +1.14% +1.45%
+Diamonds +0.60% +1.27% +0.05pp +1.01% +1.15%
+Digits +28.23% +7.48% -0.02pp +10.91% +3.09%
+Elevators -0.54% +0.13% +0.02pp +0.34% +0.67%
+House 16H -1.88% +0.42% +0.27pp +0.20% +0.61%
+House Sales -0.71% -1.35% -0.34pp +0.22% -1.02%
+Kin8nm -0.39% -0.77% -0.12pp -0.32% +0.97%
+Mercedes Benz +3.46% +1.14% +1.22pp +2.12% +1.31%
+NYC Taxi +4.71% +3.04% +3.90pp +4.19% +2.08%
+Protein -1.47% +0.51% +0.45pp +1.43% +3.34%
+Pol +7.37% +3.23% +3.64pp +8.15% +13.82%
+Puma8NH +7.24% +7.41% +0.71pp +6.26% +6.37%
+Sulfur -13.24% +0.36% +0.48pp +5.14% +18.50%
+Figure 9: Impact of CRLS fine-tuning on realTabPFNv2.5
+8.5 CRLS performances of TabICLv2 versus realTabPFNv2.5
+Table 10: Aggregate Performance Improvement (TabICLv2 vs. realTabPFNv2.5)
+Metric Improvement (mean $\pm$ std) Median W/L/T
+MAE +1.41% $\pm$ 12.65% +0.16% 56/48/1
+RMSE +1.85% $\pm$ 10.02% +0.96% 65/34/6
+$R^{2}$ +3.22pp $\pm$ 17.03pp +0.17pp 56/27/22
+CRLS +6.01% $\pm$ 10.95% +3.49% 84/19/2
+IntervalScore95 +5.14% $\pm$ 14.18% +2.15% 70/35/0
+Table 11: Dataset-Specific Improvement Metrics
+Dataset MAE Imp. RMSE Imp. $R^{2}$ Imp. CRLS Imp. IS95 Imp.
+Abalone +0.22% +0.92% +0.76pp +3.37% +4.94%
+Ailerons -1.69% -1.01% -0.29pp +1.55% +0.23%
+Ames Housing +0.37% +8.13% +32.38pp +4.95% -13.51%
+Bike Sharing +0.21% -0.62% -0.06pp +2.83% -0.13%
+Black Friday +0.59% +1.17% +1.18pp +0.34% -2.20%
+Brazilian Houses -13.82% +3.41% +6.13pp +25.02% +31.86%
+California Housing +4.87% +5.25% +1.91pp +8.33% +6.34%
+Cpu Small +1.49% +0.08% +0.00pp +6.17% +3.52%
+Diamonds +1.15% +1.62% +0.06pp +4.33% +4.93%
+Digits +42.92% +21.15% +0.89pp +34.26% +43.52%
+Elevators -0.20% +0.57% +0.10pp +1.55% -0.55%
+House 16H +0.76% +2.92% +1.92pp +5.30% +5.42%
+House Sales -5.00% -10.86% -2.37pp +0.11% -3.01%
+Kin8nm +0.59% -0.29% -0.04pp +1.52% +3.26%
+Mercedes Benz +18.83% +9.57% +9.92pp +20.31% +13.27%
+NYC Taxi -7.26% +3.99% +9.25pp -5.47% -6.48%
+Protein +2.63% +1.71% +1.55pp +4.28% +3.15%
+Pol +10.39% +4.05% +4.56pp +11.22% +7.74%
+Puma8NH -13.29% -13.86% -1.50pp -13.66% -10.89%
+Sulfur -10.13% +1.27% +1.61pp +6.38% +16.39%
+Visualizing Soil -4.08% -0.36% -0.32pp +3.50% +0.08%
+Figure 10: Impact of changing from realTabPFNv2.5 to TabICLv2: Even stronger outperformance of TabICLv2 over realTabPFNv2.5 on CRLS than that achieved by fine-tuning realTabPFNv2.5 on CRLS
+8.6 CRPS performances of realTabPFNv2.5 and TabICLv2
+Figure 11: Performance Improvement of TabICLv2 over realTabPFNv2.5 on CRPS (and other metrics)
+Table 12: Overall Model Improvement (TabICLv2 vs. realTabPFNv2.5)
+Metric Improvement (mean $\pm$ std) Median W/L/T
+MAE +1.41% $\pm$ 12.65% +0.16% 56/48/1
+RMSE +1.85% $\pm$ 10.02% +0.96% 65/34/6
+$R^{2}$ +3.22pp $\pm$ 17.03pp +0.17pp 56/27/22
+CRPS +2.01% $\pm$ 10.17% +0.73% 59/45/1
+IntervalScore95 +5.14% $\pm$ 14.18% +2.15% 70/35/0
+Table 13: Dataset-Specific Performance Gains
+Dataset MAE Imp. RMSE Imp. $R^{2}$ Imp. CRPS Imp. IS95 Imp.
+Abalone +0.22% +0.92% +0.76pp +0.86% +4.94%
+Ailerons -1.69% -1.01% -0.29pp -2.20% +0.23%
+Ames Housing +0.37% +8.13% +32.38pp -1.81% -13.51%
+Bike Sharing +0.21% -0.62% -0.06pp -0.75% -0.13%
+Black Friday +0.59% +1.17% +1.18pp +0.61% -2.20%
+Brazilian Houses -13.82% +3.41% +6.13pp +3.63% +31.86%
+California Housing +4.87% +5.25% +1.91pp +5.84% +6.34%
+Cpu Small +1.49% +0.08% +0.00pp +1.82% +3.52%
+Diamonds +1.15% +1.62% +0.06pp +1.45% +4.93%
+Digits +42.92% +21.15% +0.89pp +28.19% +43.52%
+Elevators -0.20% +0.57% +0.10pp -0.17% -0.55%
+House 16H +0.76% +2.92% +1.92pp +1.82% +5.42%
+House Sales -5.00% -10.86% -2.37pp -4.79% -3.01%
+Kin8nm +0.59% -0.29% -0.04pp +0.20% +3.26%
+Mercedes Benz +18.83% +9.57% +9.92pp +20.91% +13.27%
+NYC Taxi -7.26% +3.99% +9.25pp -8.96% -6.48%
+Protein +2.63% +1.71% +1.55pp +2.79% +3.15%
+Pol +10.39% +4.05% +4.56pp +6.73% +7.74%
+Puma8NH -13.29% -13.86% -1.50pp -15.15% -10.89%
+Sulfur -10.13% +1.27% +1.61pp +1.78% +16.39%
+Visualizing Soil -4.08% -0.36% -0.32pp -0.56% +0.08%